Jefferson Lab HALL A Logbook for April 08 - All EXCEPT Automatic Entries
Main INDEX
Wednesday
236035 04/30/08 23:14 Yi Qiang Beam Energy Strip Chart
236034 04/30/08 23:13 Yi Qiang BigBite HV Screen
236033 04/30/08 23:12 Yi Qiang Hall A Tools Screen
236030 04/30/08 22:59 Yi Qiang Target Screen
236008 04/30/08 20:07 adaq Hall A Tools snapshot
236007 04/30/08 20:07 adaq Beam Position and Current snapshot
236006 04/30/08 20:07 adaq Beam Energy in A and C snapshot
236005 04/30/08 20:01 Norum Stop data taking; beam retune
235979 04/30/08 17:55 Ramesh Carbon Hole
235966 04/30/08 16:16 Ole Hansen Day shift summary
235964 04/30/08 15:58 E.Chudakov Moller results
235963 04/30/08 15:37 tech/folts diffuser cooler fault
235962 04/30/08 15:32 brads Timelapse of balloon deflation through last fill (link)
235961 04/30/08 15:26 brads Timelapse of balloon deflation through last fill
235960 04/30/08 15:22 Ole Hansen Diffuser still faults
235959 04/30/08 15:14 ycwang Move Target from Empty to Carbon/with Hole
235958 04/30/08 15:14 Ole Hansen Resuming production
235957 04/30/08 14:39 Ole Hansen Hall status
235956 04/30/08 14:35 Higinbotham Balloons Refilled
235955 04/30/08 13:20 E.Chudakov Moller measurement
235954 04/30/08 12:07 adaq Hall A Tools snapshot
235953 04/30/08 12:07 adaq Beam Position and Current snapshot
235952 04/30/08 12:07 adaq Beam Energy in A and C snapshot
235951 04/30/08 11:33 Ramesh Energy lock at low curren
235936 04/30/08 09:12 Ole Hansen Beam back
235935 04/30/08 09:01 Ole Hansen hapc7 X display was misconfigured
235934 04/30/08 08:59 Ole Hansen Beam off for spot change
235929 04/30/08 08:35 ycwang Target Screens
235922 04/30/08 08:01 Saha Shift Summary
235914 04/30/08 07:15 A. Camsonne More about gas Cerenov PMT6 on left arm
235913 04/30/08 07:09 A. Camsonne Warning message for scaler and EPICS for right arm
235911 04/30/08 06:58 A. Camsonne Gas Cerenkov PMT 6 for run 4414
235908 04/30/08 06:54 Saha BigBite HV Trip at 0651
235900 04/30/08 06:03 Aniol front helium bag looks thin
235897 04/30/08 05:48 A. Camsonne Coincidence time normal with beam
235896 04/30/08 05:42 Saha Big Bite Hv Trip
235895 04/30/08 05:42 Saha Beam Energy Screen
235894 04/30/08 05:41 Saha BPM/BCM Screen
235893 04/30/08 05:40 Saha Hall A Tools Screen
235890 04/30/08 05:38 A. Camsonne Question about deadtime at 3.5 uA
235885 04/30/08 05:22 A. Camsonne Run 4407 coincindence time
235880 04/30/08 04:38 Aniol tgt charts
235879 04/30/08 04:07 adaq Hall A Tools snapshot
235878 04/30/08 04:07 adaq Beam Position and Current snapshot
235877 04/30/08 04:07 adaq Beam Energy in A and C snapshot
235875 04/30/08 00:13 S. Sirca Addition to the online replay GUI
Tuesday
235873 04/29/08 23:58 S. Sirca Golden run for kin C-III is 4404
235872 04/29/08 23:56 S. Sirca Swing shift summary
235870 04/29/08 23:48 S. Sirca Change in HRSL Cerenkov ADC 6
235861 04/29/08 22:45 S. Sirca Massive BB HV trip
235856 04/29/08 22:20 brads Swapped out card and power cycled LHRS HV crate
235854 04/29/08 21:12 brads LHRS HV snapshot
235853 04/29/08 21:10 S. Sirca Hall access
235852 04/29/08 20:53 brads Run 4399 has 20% true deadtime (vs. 3% from datamon) -- don't be fooled by datamon
235851 04/29/08 20:52 S. Sirca Flaky Lecroy slot (?)
235847 04/29/08 20:27 brads Computer Center portscanning our ROCs? (and what's going on with halog?)
235845 04/29/08 20:24 Armstrong Target Screen shots
235843 04/29/08 20:20 brads Computer Center scanning our ROCs?
235842 04/29/08 20:18 Computer Center portscanning our ROCs?
235841 04/29/08 20:18 Computer Center portscanning our ROCs?
235838 04/29/08 20:07 adaq Hall A Tools snapshot
235837 04/29/08 20:07 adaq Beam Position and Current snapshot
235836 04/29/08 20:06 adaq Beam Energy in A and C snapshot
235830 04/29/08 19:40 Armstrong correction re: deadtime
235822 04/29/08 19:13 Armstrong deadtime puzzle
235820 04/29/08 18:54 Armstrong move to hydrogen target
235818 04/29/08 18:37 Armstrong reduce beam current
235811 04/29/08 18:24 Armstrong moved to C target
235804 04/29/08 16:37 E.Chudakov Moller DAQ fixed
235803 04/29/08 16:33 Armstrong Target screen shots
235801 04/29/08 16:06 Ole Hansen Day shift summary
235800 04/29/08 15:05 hyao target moved from H2 to Tantalum
235799 04/29/08 13:23 Ole Hansen Kinematics C
235798 04/29/08 13:23 Higinbotham BigBite Set To 54 Degrees
235797 04/29/08 12:40 jinhuang replay software updated
235796 04/29/08 12:06 adaq Hall A Tools snapshot
235795 04/29/08 12:06 adaq Beam Position and Current snapshot
235794 04/29/08 12:06 adaq Beam Energy in A and C snapshot
235793 04/29/08 11:53 R. Michaels removing HAPPEX FLEXIO output 0
235792 04/29/08 11:33 R. Michaels HAPPEX DAQ tests
235777 04/29/08 10:06 Ole Hansen Beam studies
235769 04/29/08 09:14 Ole Hansen BB HVs back
235768 04/29/08 09:05 Ole Hansen BB MWDC HV trip
235754 04/29/08 08:07 Aniol owl shift summary
235750 04/29/08 08:04 hyao target moved to H2
235747 04/29/08 07:52 K. Aniol r4369 is 6cm Al dummy.
235744 04/29/08 07:38 Aniol 6cm Al dummy tgt in
235717 04/29/08 05:46 khem reboot all ROC
235697 04/29/08 04:06 adaq Hall A Tools snapshot
235696 04/29/08 04:06 adaq Beam Position and Current snapshot
235695 04/29/08 04:06 adaq Beam Energy in A and C snapshot
235658 04/29/08 01:39 Aniol BB HV snapshot
235653 04/29/08 01:24 Aniol beam energy strip chart
235650 04/29/08 01:20 Aniol bpm/bcm
235647 04/29/08 01:06 Adaq tgt charts
235646 04/29/08 01:04 Aniol tgt shots
235643 04/29/08 00:57 Aniol general tools screen
235638 04/29/08 00:39 brads Hall A and C momentum readbacks from EPICs have deadlined?
235635 04/29/08 00:30 adaq screen grab test
235632 04/29/08 00:25 Adaq test screen grab
235626 04/29/08 00:03 Y. Zhang swing shift summary
Monday
235621 04/28/08 23:42 Chiranjib TARGET STATUS
235606 04/28/08 22:49 Chiranjib TARGET---> L1
235604 04/28/08 22:32 Chiranjib TARGET---> C
235599 04/28/08 21:35 brads HV off on FPP now
235597 04/28/08 21:29 brads Re-enabled system beep on hacsbc2 (for alarm handler)
235578 04/28/08 20:06 adaq Hall A Tools snapshot
235577 04/28/08 20:06 adaq Beam Position and Current snapshot
235576 04/28/08 20:06 adaq Beam Energy in A and C snapshot
235567 04/28/08 19:25 Y. Zhang Hall A General Tools alarm
235552 04/28/08 17:38 Chiranjib TARGET---> L1
235547 04/28/08 16:54 E.Chudakov Can not do Moller
235546 04/28/08 16:52 Chiranjib TARGET---> Al
235541 04/28/08 16:03 jpchen day shift summary cont.
235539 04/28/08 15:59 jpchen fan to 24 Hz for loop 2
235533 04/28/08 15:29 jpchen day shift summary
235532 04/28/08 15:24 Armstrong Wein angle change
235531 04/28/08 15:24 Armstrong Wein angle change
235527 04/28/08 15:13 jpchen Big Bite chamber HVs
235526 04/28/08 15:12 jpchen Beam energy strip chart
235525 04/28/08 15:11 jpchen BPM/BCM
235524 04/28/08 15:11 jpchen tools screen
235520 04/28/08 14:56 Armstrong Another Wein angle change
235518 04/28/08 14:53 jpchen FPP test, HV settings and run #
235513 04/28/08 14:37 Higinbotham Accelerator Energies
235511 04/28/08 14:26 Armstrong Wein Angle change
235481 04/28/08 12:06 adaq Hall A Tools snapshot
235480 04/28/08 12:06 adaq Beam Position and Current snapshot
235479 04/28/08 12:06 adaq Beam Energy in A and C snapshot
235476 04/28/08 12:01 jpchen completed adjustment of coolant flow
235475 04/28/08 11:54 jpchen run 4261
235470 04/28/08 11:39 Armstrong Target screen shots
235455 04/28/08 10:09 jpchen Adjust coolant flow
235440 04/28/08 09:13 jpchen event type 5/6
235425 04/28/08 08:05 M.Shabestari Owl Shift Summary
235422 04/28/08 08:03 M.Shabestari BPM and BCM stripcharts
235413 04/28/08 07:37 M.Shabestari BigBite HV snapshot
235405 04/28/08 07:05 M.Shabestari Beam energy stripchart
235401 04/28/08 06:52 M.Shabestari Hall A Tools snapshot
235373 04/28/08 04:32 M.Shabestari Target screens
235368 04/28/08 04:20 M.Shabestari BigBite HV snapshot
235364 04/28/08 04:06 adaq Hall A Tools snapshot
235363 04/28/08 04:06 adaq Beam Position and Current snapshot
235362 04/28/08 04:06 adaq Beam Energy in A and C snapshot
235343 04/28/08 02:47 M.Shabestari Radiation stripchart
235321 04/28/08 01:15 M.Shabestari BigBite HV snapshot
235316 04/28/08 01:01 M.Shabestari Beam energy stripchart
235315 04/28/08 00:59 Aniol adaql8 replay error msg
235310 04/28/08 00:42 M.Shabestari BCM and BPM stripcharts
235309 04/28/08 00:40 M.Shabestari Hall A Tools snapshot
235301 04/28/08 00:14 Y. Zhang beam time summary disappear
235299 04/28/08 00:11 Y. Zhang Swing Shift Summary
Sunday
235296 04/27/08 23:56 Y. Zhang BB HV traped
235242 04/27/08 20:06 adaq Hall A Tools snapshot
235241 04/27/08 20:06 adaq Beam Position and Current snapshot
235240 04/27/08 20:06 adaq Beam Energy in A and C snapshot
235221 04/27/08 18:41 Y.Zhang BB HV traped
235216 04/27/08 18:18 Kalyan Target screens
235197 04/27/08 17:03 Y. Zhang BB MWDC hits plane
235186 04/27/08 16:19 Lindgren Day shift summary
235178 04/27/08 15:46 Moffit New dE thresholds for PMT singles.
235174 04/27/08 15:33 Lindgren BigBite Scaler Rates
235172 04/27/08 15:30 Lindgren Ozone/Rad strip chart
235171 04/27/08 15:29 Lindgren Ozone/Rad monitors strip chart
235170 04/27/08 15:28 Lindgren BPM/BCM Strip Chart
235169 04/27/08 15:28 Lindgren Beam Energy Strip Chart
235168 04/27/08 15:26 Lindgren BigBite HV
235163 04/27/08 15:10 Tharanga screen capture
235160 04/27/08 15:04 Lindgren Beam back on
235157 04/27/08 14:52 Tharanga target screen shot
235156 04/27/08 14:51 Tharanga target
235155 04/27/08 14:51 tharanga target
235154 04/27/08 14:50 Tharanga Target
235153 04/27/08 14:47 Lindgren Hall A Tools
235152 04/27/08 14:00 Lindgren Cavity Trip
235131 04/27/08 12:33 Lindgren Mid shift summary
235124 04/27/08 12:06 adaq Hall A Tools snapshot
235123 04/27/08 12:06 adaq Beam Position and Current snapshot
235122 04/27/08 12:06 adaq Beam Energy in A and C snapshot
235121 04/27/08 11:57 Lindgren Production Running
235103 04/27/08 11:23 Moffit Re: hapc7. HV alarm and Montecello update
235092 04/27/08 10:52 Lindgren Setting up for kinematics L
235082 04/27/08 10:27 Moffit Swapped out F1TDC
235077 04/27/08 10:21 D McNulty Move to Ta target and prepare for Kinematic L
235076 04/27/08 10:11 Moffit Reset BB TP thresholds
235074 04/27/08 10:06 techs left spectrometer to 12.5
235063 04/27/08 09:29 Moffit Reduced BB tp thresholds for cosmics
235062 04/27/08 09:17 R.Lindgren Hall A Access
235051 04/27/08 08:10 M.Shabestari Owl Shift Summary
235049 04/27/08 08:00 M.Shabestari Target screens
235042 04/27/08 07:35 M.Shabestari Beam energy stripchart
235041 04/27/08 07:32 M.Shabestari BigBite HV snapshot
235024 04/27/08 06:34 M.Shabestari Hall A Tools snapshot
235005 04/27/08 04:53 M.Shabestari BCM and BPM stripcharts
234997 04/27/08 04:35 M.Shabestari BigBite HV snapshot
234996 04/27/08 04:32 M.Shabestari Radiation stripchart
234990 04/27/08 04:06 adaq Hall A Tools snapshot
234989 04/27/08 04:06 adaq Beam Position and Current snapshot
234988 04/27/08 04:06 adaq Beam Energy in A and C snapshot
234985 04/27/08 03:59 M.Shabestari Target screens
234961 04/27/08 02:15 M.Shabestari HV tripped due to beam trip
234957 04/27/08 01:55 M.Shabestari BigBite HV snapshot
234948 04/27/08 01:15 M.Shabestari Beam energy stripchart
234946 04/27/08 01:09 M.Shabestari HV tripped
234942 04/27/08 00:56 M.Shabestari BCM and BPM stripcharts
234939 04/27/08 00:43 M.Shabestari Hall A Tools snapshot
234934 04/27/08 00:31 M.Shabestari Target screens
234929 04/27/08 00:16 Ramesh Shift summary
Saturday
234913 04/26/08 23:00 Ramesh BPM and BCM stripcharts
234912 04/26/08 22:58 Ramesh BB HV gui
234911 04/26/08 22:56 Ramesh Energy stripchart, BB HV gui
234910 04/26/08 22:54 Ramesh Hall A tools, beam current monitor
234904 04/26/08 22:35 Chiranjib TARGET----->> L1
234903 04/26/08 22:31 Ramesh Al dummy runs finished
234894 04/26/08 21:22 Ramesh Al dummy run started
234893 04/26/08 21:15 Chiranjib TARGET--->>Al
234891 04/26/08 20:58 R. Michaels re: helicity errors
234886 04/26/08 20:41 Chiranjib Target status
234879 04/26/08 20:14 Ramesh BB HV trip, now OK
234877 04/26/08 20:06 adaq Hall A Tools snapshot
234876 04/26/08 20:06 adaq Beam Position and Current snapshot
234875 04/26/08 20:06 adaq Beam Energy in A and C snapshot
234874 04/26/08 20:04 Chiranjib WARNING THaMyG0Helicity
234873 04/26/08 19:52 Chiranjib Helicity check for RUN 4029
234872 04/26/08 19:50 Ramesh S5 HV
234871 04/26/08 19:08 chiranjib RUN 4025 helicity
234852 04/26/08 17:32 Ramesh BB HV trip (late entry)
234849 04/26/08 17:25 Chiranjib Target status
234846 04/26/08 17:21 Chiranjib current helicity check
234841 04/26/08 16:16 Lindgren BigBite Scalers
234840 04/26/08 16:14 Lindgren BPM/BCM
234837 04/26/08 16:10 Lindgren Beam current and energy plots
234836 04/26/08 16:08 Lindgren Hall A Tools
234835 04/26/08 16:07 Lindgren Day Shift summry
234804 04/26/08 13:48 Aniolo target charts
234785 04/26/08 12:15 R. Lindgren Mid-shift summary
234782 04/26/08 12:06 adaq Hall A Tools snapshot
234781 04/26/08 12:06 adaq Beam Position and Current snapshot
234780 04/26/08 12:06 adaq Beam Energy in A and C snapshot
234753 04/26/08 09:59 Lindgren HV crashed
234748 04/26/08 09:41 Lindgren Update
234741 04/26/08 09:10 Lindgren Day shift begining
234728 04/26/08 08:10 M.Shabestari Target screens
234726 04/26/08 08:09 M.Shabestari Owl Shift Summary
234717 04/26/08 07:35 khem right arm end 20768 and start 20769 are mess up
234700 04/26/08 06:23 M.Shabestari BPM and BCM stripcharts
234699 04/26/08 06:19 M.Shabestari BigBite HV snapshot
234697 04/26/08 06:18 M.Shabestari Hall A Tools snapshot
234665 04/26/08 04:06 adaq Hall A Tools snapshot
234664 04/26/08 04:05 adaq Beam Position and Current snapshot
234663 04/26/08 04:05 adaq Beam Energy in A and C snapshot
234662 04/26/08 04:05 M.Shabestari Target screens
234661 04/26/08 04:01 M.Shabestari Radiation stripchart
234638 04/26/08 02:40 M.Shabestari HVs snapsot
234637 04/26/08 02:37 M.Shabestari HV lost connection
234632 04/26/08 02:23 M.Shabestari Beam energy stripchart
234631 04/26/08 02:20 M.Shabestari Target screens & notes
234624 04/26/08 01:53 M.Shabestari Target: Fan speed
234622 04/26/08 01:52 M.Shabestari Target: ESR_Flow
234616 04/26/08 01:30 M.Shabestari BigBite HV snapshot
234611 04/26/08 01:16 M.Shabestari Hall A Tools snapshot
234610 04/26/08 01:15 M.Shabestari BCM and BPM snapshots
234602 04/26/08 00:44 M.Shabestari HV lost connection
234593 04/26/08 00:09 A.Rakhman Shift Summary Swing Shift
234591 04/26/08 00:04 A.Rakhman Beam Trip
Friday
234584 04/25/08 23:45 A.Rakhman BB HV Slot 14 Screen Shot
234583 04/25/08 23:44 A.Rakhman Beam Energy Strip Chart Screen Shot
234582 04/25/08 23:43 Y. Zhang BB MWDC hits per plane
234581 04/25/08 23:42 A.Rakhman BCM BPM Strip Chart Screen Shot
234580 04/25/08 23:40 A.Rakhman General Tools Screen Shot
234577 04/25/08 23:37 Moffit Changed golden run to 3864
234555 04/25/08 22:06 A.Rakhman BigBite HV Slot 14
234554 04/25/08 22:04 A.Rakhman Beam Energy Strip Chart
234553 04/25/08 22:03 A.Rakhman BPM & BCM Strip Chart
234552 04/25/08 22:02 Y.Zhang Target snapshot
234551 04/25/08 22:02 A.Rakhman Hall A General Tools
234546 04/25/08 21:46 A.Rakhman Snap Shot
234545 04/25/08 21:45 A.Rakhman Snap Shot
234536 04/25/08 21:25 Y Zhang BB MWDC hits & HRS scatter angle
234535 04/25/08 21:22 Y.Zhang BB MWDC hits & HRS scatter angle
234522 04/25/08 20:27 A.Rakhman Snap Shot
234521 04/25/08 20:25 A.Rakhman Snap Shot
234514 04/25/08 20:05 adaq Hall A Tools snapshot
234513 04/25/08 20:05 adaq Beam Position and Current snapshot
234512 04/25/08 20:05 adaq Beam Energy in A and C snapshot
234463 04/25/08 16:32 A.Rakhman Snap Shot 2
234462 04/25/08 16:30 A.Rakhman Snap Shots
234455 04/25/08 16:08 Lindgren Shift Summary
234444 04/25/08 15:31 Lindgren Current
234440 04/25/08 15:13 mercado Target Charts
234430 04/25/08 14:35 R. Michaels simplifed helicity signal path
234427 04/25/08 14:32 Lindgren beam current increase again
234419 04/25/08 14:10 R. Michaels helicity crisis, recommended actions
234414 04/25/08 13:54 Lindgren MWDC HV
234411 04/25/08 13:45 Lindgren Ozone/Rad Mon
234410 04/25/08 13:44 Lindgren Hall A Tools
234409 04/25/08 13:43 Lindgren Beam Chart
234407 04/25/08 13:38 Lindgren beam increase
234405 04/25/08 13:35 Moffit Set golden run to 3835.
234404 04/25/08 13:28 Lindgren Mid Day Summary
234389 04/25/08 12:22 Lindgren Prescales
234384 04/25/08 12:05 adaq Hall A Tools snapshot
234383 04/25/08 12:05 adaq Beam Position and Current snapshot
234382 04/25/08 12:05 adaq Beam Energy in A and C snapshot
234368 04/25/08 10:41 Lindgren Production Running
234366 04/25/08 10:40 Lindgren Changed probe
234360 04/25/08 10:08 Lindgren Changing to LH2 target
234359 04/25/08 10:05 Lindgren Carbon target
234357 04/25/08 09:54 V. Sulkosky Helicity for Run 3462
234352 04/25/08 09:22 Lindgren Beam current
234347 04/25/08 09:13 Lindgren lower beam current
234345 04/25/08 09:11 R. Lindgren Angle change
234343 04/25/08 08:32 Moffit RADCON and replacing F1TDC
234336 04/25/08 08:08 ycwang shiftsummary (owl) 04/25/08
234334 04/25/08 08:05 youcai wang shift summary (owl) 04/25/08
234333 04/25/08 08:05 youcai wang shift summary (owl) 04/25/08
234332 04/25/08 08:02 Youcai Wang Shift summary (owl shift 04/25/08)
234330 04/25/08 08:02 Youcai Wang Shift summary (owl shift 04/25/08)
234320 04/25/08 07:18 D. McNulty 4/24/2008 DAY SHIFT SUMMARY
234289 04/25/08 04:49 ycwang BigBite scalers
234288 04/25/08 04:48 ycwang BPM/BCM
234287 04/25/08 04:47 ycwang Compton/OTR
234286 04/25/08 04:46 ycwang Beam Energy
234285 04/25/08 04:46 ycwang Ozone/Rad Mon
234284 04/25/08 04:45 ycwang BigBite HV
234283 04/25/08 04:44 ycwang HV Right/Left
234282 04/25/08 04:44 ycwang Hall A General Tools
234271 04/25/08 04:05 adaq Hall A Tools snapshot
234270 04/25/08 04:05 adaq Beam Position and Current snapshot
234269 04/25/08 04:05 adaq Beam Energy in A and C snapshot
234244 04/25/08 02:10 khem no roc5 warning message any more
234233 04/25/08 01:09 khem roc 5 warning message
234226 04/25/08 00:46 ycwang BigBite Scalers
234225 04/25/08 00:46 ycwang BPM/BCM
234224 04/25/08 00:45 ycwang Compton/OTR
234220 04/25/08 00:43 ycwang Beam Energy
234218 04/25/08 00:43 ycwang Ozone/ Rad Mon
234217 04/25/08 00:42 ycwang Bigbite HV
234216 04/25/08 00:41 ycwang Hall A General Tools
234215 04/25/08 00:40 ycwang HV Right/Left
234211 04/25/08 00:19 ycwang target screens
234206 04/25/08 00:03 LeRose Shift Summary (swing)
Thursday
234195 04/24/08 23:08 LeRose ROC5
234175 04/24/08 20:53 mercado Run 3748?
234174 04/24/08 20:49 brads He bag not to blame for 3732
234169 04/24/08 20:41 LeRose BPM/BCM BB HV
234168 04/24/08 20:37 LeRose Hall A Tools/Energy
234161 04/24/08 20:05 adaq Hall A Tools snapshot
234160 04/24/08 20:05 adaq Beam Position and Current snapshot
234159 04/24/08 20:05 adaq Beam Energy in A and C snapshot
234146 04/24/08 18:45 mercado Target Charts
234129 04/24/08 17:31 mercado Filling up the Helium Bag
234120 04/24/08 16:46 D. McNulty Is parity DAQ affecting Counting DAQ helicity?
234105 04/24/08 15:29 Norum Target status
234091 04/24/08 14:04 Saha BPM/BCM Screen
234089 04/24/08 14:01 Saha Beam Energy screen
234088 04/24/08 14:00 Saha Hall A Tools
234081 04/24/08 13:25 brads BB He balloon needs refilling
234080 04/24/08 13:24 Moffit Changed golden run to 3714
234065 04/24/08 12:05 adaq Hall A Tools snapshot
234064 04/24/08 12:05 adaq Beam Position and Current snapshot
234063 04/24/08 12:05 adaq Beam Energy in A and C snapshot
234058 04/24/08 10:09 brads TDC window change from yesterday looks OK
234039 04/24/08 08:09 B. Hahn Owl Shift Summary
234033 04/24/08 07:56 R. Michaels 0.106 Coul so far (Kine J)
234008 04/24/08 05:24 R. Michaels putting ps6 to 1
233997 04/24/08 04:24 R. Michaels ps6 has been 2
233996 04/24/08 04:22 B. Hahn Beam Energy in Hall A and C
233995 04/24/08 04:21 B. Hahn Hall A Tools and Beam Position and Current
233992 04/24/08 04:19 B. Hahn Big Bite HV snapshot
233988 04/24/08 04:05 adaq Hall A Tools snapshot
233987 04/24/08 04:05 adaq Beam Position and Current snapshot
233986 04/24/08 04:05 adaq Beam Energy in A and C snapshot
233985 04/24/08 04:03 R. Michaels target screens (II)
233984 04/24/08 04:02 R. Michaels target screens (I)
233980 04/24/08 03:45 R. Michaels hapc7 rebuilt
233964 04/24/08 02:10 R. Michaels ps4 to 55
233959 04/24/08 02:02 B. Hahn Very few T4 events
233956 04/24/08 01:47 Moffit Re: golden run comparisons..
233951 04/24/08 01:18 B. Hahn Differences from Golden Run in online analysis
233938 04/24/08 00:26 R. Michaels moving to production
233932 04/24/08 00:00 R. Michaels going to Hyd. elastic, Kine J
Wednesday
233929 04/23/08 23:55 B. Norum Status at shift end
233928 04/23/08 23:50 Yi Qiang Hall A Tools Screen
233921 04/23/08 22:58 B Norum Beam restore
233920 04/23/08 20:05 adaq Hall A Tools snapshot
233919 04/23/08 20:05 adaq Beam Position and Current snapshot
233918 04/23/08 20:05 adaq Beam Energy in A and C snapshot
233917 04/23/08 20:00 V. Sulkosky Golden Rootfile change to 3670.
233916 04/23/08 19:57 V. Sulkosky New Helicity Historgram in Replay
233915 04/23/08 16:38 khem set T1 Threshold to 550
233913 04/23/08 15:58 Ole Hansen Day Shift Summary
233912 04/23/08 15:58 Khem T1 trigger threshold
233910 04/23/08 13:54 Ole Hansen Still no beam. ETA 1 hour.
233909 04/23/08 13:21 Moffit Re: window change... of course.. it's 1000 nanoseconds.
233908 04/23/08 13:08 Moffit Changed MWDC TDC window to 1000 microseconds
233907 04/23/08 12:05 adaq Hall A Tools snapshot
233906 04/23/08 12:05 adaq Beam Position and Current snapshot
233905 04/23/08 12:05 adaq Beam Energy in A and C snapshot
233904 04/23/08 09:38 Ole Hansen Going to kinematics J
233903 04/23/08 08:08 M.Shabestari Owl Shift Summary
233901 04/23/08 08:05 M.Shabestari Target screens
233900 04/23/08 07:10 M.Shabestari Hall A Tools snapshot
233899 04/23/08 07:09 M.Shabestari BCM and BPM snapshots
233886 04/23/08 04:44 M.Shabestari Target screens
233881 04/23/08 04:15 M.Shabestari BigBite HV snapshot
233880 04/23/08 04:05 adaq Hall A Tools snapshot
233879 04/23/08 04:05 adaq Beam Position and Current snapshot
233878 04/23/08 04:05 adaq Beam Energy in A and C snapshot
233875 04/23/08 03:41 M.Shabestari Radiation stripchart
233874 04/23/08 03:39 M.Shabestari Beam energy stripchart
233865 04/23/08 02:51 M.Shabestari Timing comparison
233862 04/23/08 02:27 Moffit APC log from hareboot30
233859 04/23/08 02:14 Moffit Timing appears to be restored
233856 04/23/08 01:55 M.Shabestari BCM and BPM snapshots
233855 04/23/08 01:51 khem T3 T4 T7 timing run# 3629,3655
233854 04/23/08 01:47 M.Shabestari ROC5 harebooted
233853 04/23/08 01:44 khem hareboot roc5
233852 04/23/08 01:44 M.Shabestari Hall A Tools snapshot
233849 04/23/08 01:25 M.Shabestari Target screens
233848 04/23/08 01:19 M.Shabestari HV trip
233845 04/23/08 01:02 khem controlled access to reboot roc9,10
233844 04/23/08 01:00 M.Shabestari Threshold on MWDC
233838 04/23/08 00:27 Yi Qiang Swing Shift Summary
233836 04/23/08 00:20 Y. Zhang BPM & Compton snapshot
233833 04/23/08 00:18 Y. Zhang Hall A general snapshot
233832 04/23/08 00:04 Yi Qiang Controlled Access to Reboot ROC9/10
Tuesday
233831 04/22/08 23:52 Moffit BigBite Camac Controller was reset.
233830 04/22/08 23:27 Y. Zhang target snapshot
233824 04/22/08 21:31 E.Chudakov Charge asymmetry with Moller
233821 04/22/08 20:56 Yi Qiang Beam Trips
233820 04/22/08 20:42 Yi Qiang A detailed description of target after the power trip
233817 04/22/08 20:30 Yi Qiang Loop1 HPH recovered itself
233816 04/22/08 20:28 Yi Qiang Target -> Al
233815 04/22/08 20:23 Yi Qiang Target High Power Heater Gone Crazy due to Another Power Trip
233812 04/22/08 20:05 adaq Hall A Tools snapshot
233811 04/22/08 20:04 adaq Beam Position and Current snapshot
233810 04/22/08 20:04 adaq Beam Energy in A and C snapshot
233801 04/22/08 19:00 E.Chudakov Time structure of helicity signals
233792 04/22/08 17:46 Ole Hansen Re: track multiplicities
233778 04/22/08 16:00 Adaq run summary
233771 04/22/08 14:53 Tharanga BPM & BCM
233770 04/22/08 14:52 Tharanga Tool and BPM
233769 04/22/08 14:46 K Wang beam's back
233762 04/22/08 12:15 brads BB magnet ramped to 700A for field meas., now back to 390A
233761 04/22/08 12:04 adaq Hall A Tools snapshot
233760 04/22/08 12:04 adaq Beam Position and Current snapshot
233759 04/22/08 12:04 adaq Beam Energy in A and C snapshot
233758 04/22/08 10:53 D. McNulty parity daq charge asymmetry
233757 04/22/08 09:59 V. Sulkosky BigBite Angle Measurement
233753 04/22/08 08:19 M.Shabestari Owl Shift Summary
233752 04/22/08 08:19 K Wang HV tripped
233750 04/22/08 08:11 M.Shabestari Target screens
233748 04/22/08 08:05 M.Shabestari Hall A Tools snapshot
233745 04/22/08 07:46 M.Shabestari BigBite HV snapshot
233744 04/22/08 07:45 M.Shabestari Beam energy stripchart
233733 04/22/08 06:43 V. Sulkosky QRT Plot from 3462
233728 04/22/08 06:04 M.Shabestari BCM and BPM stripcharts
233721 04/22/08 05:08 M.Shabestari Incorrect "End of Run" for Entry number 233717
233710 04/22/08 04:04 adaq Hall A Tools snapshot
233709 04/22/08 04:04 adaq Beam Position and Current snapshot
233708 04/22/08 04:04 adaq Beam Energy in A and C snapshot
233707 04/22/08 04:01 M.Shabestari Target screens
233700 04/22/08 03:03 M.Shabestari Radiation Stripchart
233699 04/22/08 03:01 M.Shabestari BigBite HV snapshot
233697 04/22/08 02:50 M.Shabestari HV Tripped: S7 & S14
233695 04/22/08 02:41 M.Shabestari Beam energy stripchart
233694 04/22/08 02:40 M.Shabestari ROC5 rebooted
233686 04/22/08 01:41 M.Shabestari BigBite HV snapshot
233683 04/22/08 01:27 M.Shabestari Hall A Tools snapshot
233680 04/22/08 01:08 M.Shabestari Target screens
233679 04/22/08 01:05 M.Shabestari BCM and BPM stripcharts
233676 04/22/08 00:58 M.Shabestari Switched golden run to 3596 for kinematic I
233671 04/22/08 00:09 V. Sulkosky Re: Moller - QRT signal is inverted
233669 04/22/08 00:04 E.Chudakov Swing shift summary
Monday
233657 04/21/08 22:35 jinhuang Helicity support in Physcs replay
233649 04/21/08 21:37 E.Chudakov Moller - QRT signal is inverted
233640 04/21/08 20:20 R. Michaels re: RHRS Qasy
233637 04/21/08 20:04 adaq Hall A Tools snapshot
233636 04/21/08 20:04 adaq Beam Position and Current snapshot
233635 04/21/08 20:04 adaq Beam Energy in A and C snapshot
233634 04/21/08 20:04 R. Lind Moved BigBite to 43.5straight
233633 04/21/08 19:41 E.Chudakov New kinematics
233630 04/21/08 18:58 V. Sulkosky R-HRS Charge Asymmetry?
233629 04/21/08 18:49 jinhuang Physics Online Plots Updated
233628 04/21/08 18:48 jinhuang Online replay software updated
233625 04/21/08 18:22 V. Sulkosky R-HRS Large Aq Runs
233624 04/21/08 17:43 V. Sulkosky Induced Large Charge Asymmetry for Helicity and Sign Checks
233621 04/21/08 17:03 Track multiplicities
233614 04/21/08 16:18 B. Norum ROC 4 failure
233611 04/21/08 16:02 jpchen Day shift summary
233604 04/21/08 15:29 jpchen beam back, 5 uA
233603 04/21/08 15:03 V. Sulkosky Charge Asymmetry runs 3521 and 3462
233602 04/21/08 15:01 jpchen Beam off, RF problem
233599 04/21/08 14:35 jpchen track numbers @ 3.9 uA
233596 04/21/08 14:29 V. Sulkosky Charge Asymmetry Runs 3515 and 3520
233593 04/21/08 14:02 jpchen beam current to 4 uA
233590 04/21/08 13:54 jpchen BigBite chmaber HVs
233589 04/21/08 13:53 jpchen Energy strip chart
233588 04/21/08 13:52 jpchen Hall A Tools screen
233587 04/21/08 13:51 jpchen BPM/BCM strip chart
233586 04/21/08 13:48 jpchen target screens
233583 04/21/08 13:33 jpchen track numbers with 5.4 uA
233574 04/21/08 12:21 jpchen beam current reduced to 5.5 uA
233571 04/21/08 12:04 adaq Hall A Tools snapshot
233570 04/21/08 12:04 adaq Beam Position and Current snapshot
233569 04/21/08 12:04 adaq Beam Energy in A and C snapshot
233566 04/21/08 11:51 jpchen number of tracks with increasing beam current
233563 04/21/08 11:35 jpchen raster size
233561 04/21/08 11:23 jpchen raster size
233553 04/21/08 10:58 V. Sulkosky Run 3569 is Junk
233552 04/21/08 10:51 M. Mihovilov ROC 3 problem
233544 04/21/08 09:41 jpchen Beam shot up in trip recovering
233541 04/21/08 09:18 R. Michaels possibly missing MPS on R-HRS
233540 04/21/08 09:13 jpchen run 3565 and chamber 2
233537 04/21/08 08:53 jpchen ROC 5 trip
233536 04/21/08 08:52 jpchen Chmaber HV trip
233532 04/21/08 08:28 Lindgren Accumulated charge
233530 04/21/08 08:25 R. Lindgren Day shift beginning
233528 04/21/08 08:12 M.Shabestari Owl Shift Summary
233527 04/21/08 08:10 M.Shabestari Target Screens
233522 04/21/08 07:46 M.Shabestari Hall A Tools snapshot
233521 04/21/08 07:46 M.Shabestari Beam energy stripchart
233520 04/21/08 07:45 M.Shabestari BCM and BPM snapshots
233503 04/21/08 04:55 M.Shabestari Hall A Current Monitor
233500 04/21/08 04:48 M.Shabestari BigBite HV snapshot
233499 04/21/08 04:36 M.Shabestari Target screens
233496 04/21/08 04:04 adaq Hall A Tools snapshot
233495 04/21/08 04:04 adaq Beam Position and Current snapshot
233494 04/21/08 04:04 adaq Beam Energy in A and C snapshot
233481 04/21/08 01:59 M.Shabestari Radiation Stripchart
233480 04/21/08 01:57 M.Shabesatri Beam energy stripchart
233479 04/21/08 01:56 M.Shabestari BCM & BPM snapshots
233472 04/21/08 01:18 M.Shabestari Target screens
233471 04/21/08 01:14 M.Shabestari BigBite HV snapshot
233469 04/21/08 01:01 M.Shabestari HV trip on S1, S2, S3, S5, S6, S14
233467 04/21/08 00:54 M.Shabestari Hall A Tools snapshot
233466 04/21/08 00:53 M.Shabestari BPM and BCM snapshots
233463 04/21/08 00:11 Kalyan Fan on loop1 heat load
233459 04/21/08 00:08 Kalyan loop 1 fan goes to zero sometimes
233456 04/21/08 00:03 Ramesh Subed Shift summary
Sunday
233455 04/20/08 23:57 Ramesh Trippy beam
233450 04/20/08 22:53 Ramesh BPM/BCM, BigBite HV
233447 04/20/08 22:50 Ramesh Beam energy strip chart
233446 04/20/08 22:47 Ramesh Hall A tools, and BB power supply
233440 04/20/08 21:46 Ramesh BB wire chamber HV
233438 04/20/08 21:36 Ramesh ROC5 failure due to BB HV trip
233431 04/20/08 21:18 Ramesh ROC5 crashed
233418 04/20/08 20:04 adaq Hall A Tools snapshot
233417 04/20/08 20:04 adaq Beam Position and Current snapshot
233416 04/20/08 20:04 adaq Beam Energy in A and C snapshot
233404 04/20/08 18:23 V. Sulkosky Helicity and Charge Asymmetry
233402 04/20/08 18:01 Kalyan target screens
233396 04/20/08 17:34 Lindgren Shift summary
233395 04/20/08 17:27 R. Lindgren Balloon and wire chamber check
233391 04/20/08 16:17 Ramesh Energy Strip chart (late entry)
233383 04/20/08 15:23 miham Target Screens
233377 04/20/08 14:32 V. Sulkosky Golden Rootfile Changed to 3515
233376 04/20/08 14:22 V. Sulkosky Scalers in Replay
233373 04/20/08 13:45 V. Sulkosky Scalers for Run 3462
233368 04/20/08 13:18 Lindgren Momentum change for production running
233359 04/20/08 12:49 Lindgren checked Bigbite Current - ok
233358 04/20/08 12:46 Lindgren Move to LH2 target
233353 04/20/08 12:12 Lindgren Carbon target
233352 04/20/08 12:04 adaq Hall A Tools snapshot
233351 04/20/08 12:04 adaq Beam Position and Current snapshot
233350 04/20/08 12:04 adaq Beam Energy in A and C snapshot
233347 04/20/08 11:43 Lindgren Ta Calibration
233346 04/20/08 11:27 V. Sulkosky EPICS Part II and Energy variables.
233343 04/20/08 11:20 Lindgren kinematics H
233342 04/20/08 10:59 V. Sulkosky EPICS Information in Data
233325 04/20/08 09:02 R.Lindgren Collimator
233320 04/20/08 08:26 R. Michaels re: scalers and EPICS
233317 04/20/08 08:12 A. Camsonne Owl shift summary
233315 04/20/08 08:04 M.Shabestari Target screens
233312 04/20/08 08:01 A. Camsonne Scaler and EPICS in decoded runs
233311 04/20/08 07:58 A. Camsonne Helicity for run 3462 higher statistics
233302 04/20/08 06:49 A. Camsonne BPM / BCM stripchart
233301 04/20/08 06:48 A. Camsonne Hall A tools
233300 04/20/08 06:47 A. Camsonne Beam energy strip chart
233298 04/20/08 06:46 A. Camsonne Radiation stripchart
233295 04/20/08 06:36 A. Camsonne BB HV for MWDC at 4.7 uA LH2 with collimator
233294 04/20/08 06:33 M.Shabestari Target moved to LH2+Collimator
233282 04/20/08 05:28 M.Shabestari Target screens
233267 04/20/08 04:04 adaq Hall A Tools snapshot
233266 04/20/08 04:04 adaq Beam Position and Current snapshot
233265 04/20/08 04:04 adaq Beam Energy in A and C snapshot
233242 04/20/08 01:52 A. Camsonne BB scalers 4.7 uA LH2
233239 04/20/08 01:39 A. Camsonne Funny helicity for run 3462
233224 04/20/08 00:24 M.Shabestari Target screens
Saturday
233217 04/19/08 23:59 D. McNulty Swing Shift Summary: 4/19/2008
233212 04/19/08 23:40 D. McNulty MCC called about energy drift problem
233209 04/19/08 23:22 Chiranjib Target Screens
233200 04/19/08 22:34 Moffit Re: ROC5 Warnings.
233195 04/19/08 22:09 D. McNulty Screen Snapshots: Beam Energy Stripchart, BB HV slot14 (loop1 LH2, 4.8uA)
233194 04/19/08 22:05 D. McNulty Screen Snapshots: Hall A tools, BPM/BCM stripcharts
233180 04/19/08 21:04 D. McNulty Called MCC about energy drift problem
233177 04/19/08 20:54 Moffit ROC5 warnings. crl changed
233176 04/19/08 20:46 D. McNulty Energy drifted up noticeably
233171 04/19/08 20:19 D. McNulty ROC 5 died during run 3442
233170 04/19/08 20:17 Chiranjib Target screens
233168 04/19/08 20:04 adaq Hall A Tools snapshot
233167 04/19/08 20:04 adaq Beam Position and Current snapshot
233166 04/19/08 20:04 adaq Beam Energy in A and C snapshot
233141 04/19/08 17:28 Chiranjib TARGET SCREENs I
233132 04/19/08 16:39 D. McNulty TS11 crashed
233130 04/19/08 16:26 Lindgren moffits kinematic summary
233126 04/19/08 16:03 Lindgren Shift Summary
233117 04/19/08 15:18 Lindgren Target charts
233111 04/19/08 14:37 Lindgren FSD Trip
233107 04/19/08 13:51 Lindgren Accelerator Screen
233104 04/19/08 13:48 Lindgren Ozone/Rad Mon
233103 04/19/08 13:47 Lindgren Beam Energy
233102 04/19/08 13:45 Lindgren BigBite HV
233101 04/19/08 13:42 Lindgren Beam Current
233086 04/19/08 12:04 adaq Hall A Tools snapshot
233085 04/19/08 12:04 adaq Beam Position and Current snapshot
233084 04/19/08 12:04 adaq Beam Energy in A and C snapshot
233083 04/19/08 12:03 Lindgren FSD Trip
233072 04/19/08 11:05 R.Lindgren Deadtime high
233069 04/19/08 10:48 Lindgren Scalers
233068 04/19/08 10:46 Lindgren Hall A Tools and BCM
233067 04/19/08 10:38 Lindgren Running
233054 04/19/08 09:18 R.Lindgren Day shift starting
233047 04/19/08 08:41 Lindgren Beam energy strip chart
233042 04/19/08 08:08 A. Camsonne Compton strip chart
233039 04/19/08 08:05 A. Camsonne Owl shift summary
233037 04/19/08 07:59 vladimir LH2 target snapshot
233036 04/19/08 07:54 vladimir LH2 6 cm target screen
233034 04/19/08 07:45 A. Camsonne Dummy target run plots
233031 04/19/08 07:21 vladimir 6 cm dummy target screen
233029 04/19/08 07:04 vladimir target screen
233026 04/19/08 06:56 A. Camsonne Radiation stripchart
233024 04/19/08 06:44 A. Camsonne BB MWDC HV at 4.2 uA 6 cm LH2
233022 04/19/08 06:33 A. Camsonne ROC5 warning - High deadtime run 3386
232997 04/19/08 04:03 adaq Hall A Tools snapshot
232996 04/19/08 04:03 adaq Beam Position and Current snapshot
232995 04/19/08 04:03 adaq Beam Energy in A and C snapshot
232984 04/19/08 02:53 A. Camsonne Run 3368 high deadtime - ROC 5
232981 04/19/08 02:34 A. Camsonne BB wire chamber HV
232974 04/19/08 01:41 A. Camsonne Previous reference run was at lower current - Updated reference to run 3365
232967 04/19/08 01:10 A. Camsonne No scaler for run 3363
232963 04/19/08 00:54 A. Camsonne Slightly higher number of hit in BigBite compared to reference run
232957 04/19/08 00:13 vladimir Target 6 cm LH2 screen
232953 04/19/08 00:00 S. Sirca Swing shift summary
Friday
232945 04/18/08 23:27 Seonho Choi Target Screen Capture
232944 04/18/08 23:26 Seonho Choi Target Screen Capture
232943 04/18/08 23:26 Seonho Choi Target Screen Capture
232936 04/18/08 22:48 V. Sulkosky ELOG Entries on Hall A Dipole Adjustments
232933 04/18/08 22:30 S. Sirca Raster x vs. y for 12C target w/ hole
232927 04/18/08 22:12 S. Sirca Jump in beam energy, but less drift
232913 04/18/08 21:13 S. Sirca Harp scan results (H03B only)
232912 04/18/08 20:03 adaq Hall A Tools snapshot
232911 04/18/08 20:03 adaq Beam Position and Current snapshot
232910 04/18/08 20:03 adaq Beam Energy in A and C snapshot
232908 04/18/08 18:27 jinhuang Identifying Pi0 Missing Mass Peak
232907 04/18/08 17:49 Moffit New BigBite Trigger Plane scaler GUI.
232905 04/18/08 16:44 V. Sulkosky R-HRS HVs Updated
232904 04/18/08 16:08 bogdanw shift summary
232902 04/18/08 15:59 bogdanw HV setting of BB MWDC
232901 04/18/08 15:14 Moffit Relocated HV audible alarms to hapc5
232899 04/18/08 14:59 bogdanw HRS-R Hv screens
232898 04/18/08 14:55 bogdanw BB and HRS-L HV screen
232897 04/18/08 14:38 Ole Hansen MCC EPICS screens redirected to adaql1
232896 04/18/08 14:20 Ole Hansen hapc7 (MCC EPICS display) down
232895 04/18/08 14:16 Ole Hansen Still no beam
232892 04/18/08 12:30 Adaq MSS update
232891 04/18/08 12:03 adaq Hall A Tools snapshot
232890 04/18/08 12:03 adaq Beam Position and Current snapshot
232889 04/18/08 12:03 adaq Beam Energy in A and C snapshot
232888 04/18/08 11:41 LIndgren Helium cylinders
232887 04/18/08 11:36 Lindgren R-HRS filed set correctly
232886 04/18/08 11:36 Lindgren R-HRS filed set correctly
232885 04/18/08 11:33 Lindgren Helium Fill
232884 04/18/08 10:37 Moffit Changed EDTM rate to 10Hz
232883 04/18/08 10:21 Adag HRS-R momentum
232882 04/18/08 10:13 Moffit Changed nped for ROC5
232880 04/18/08 10:12 Nilanga Shift Summary
232861 04/18/08 07:22 vladimir target screen
232836 04/18/08 04:03 adaq Hall A Tools snapshot
232835 04/18/08 04:03 adaq Beam Position and Current snapshot
232834 04/18/08 04:03 adaq Beam Energy in A and C snapshot
232831 04/18/08 03:47 vladimir LH2 target snapshot
232809 04/18/08 00:07 brads Beam energy drifting again...
232808 04/18/08 00:05 S. Sirca Swing shift summary
232805 04/18/08 00:00 S. Sirca Beam energy stripchart
232803 04/18/08 00:00 S. Sirca BB HV Slot14 snapshot
Thursday
232802 04/17/08 23:57 S. Sirca BPM/BCM screens
232801 04/17/08 23:56 S. Sirca Hall A General Tools
232786 04/17/08 22:01 adaq Hall A Tools snapshot
232785 04/17/08 22:01 adaq Beam Position and Current snapshot
232784 04/17/08 22:00 adaq Beam Energy in A and C snapshot
232761 04/17/08 19:08 adaq Hall A Tools snapshot
232760 04/17/08 19:08 adaq Beam Position and Current snapshot
232759 04/17/08 19:08 adaq Beam Energy in A and C snapshot
232755 04/17/08 18:42 V. Sulkosky Beam Energy Drifting
232743 04/17/08 17:35 Moffit Switched golden run to 3274 for kinematic G
232731 04/17/08 16:01 R. Michaels day shift summary
232726 04/17/08 15:30 R. Michaels energy lock to hall A
232720 04/17/08 14:47 R. Michaels BB HV trip / recovery
232718 04/17/08 14:39 R. Michaels re: online deadtime.
232713 04/17/08 13:57 Norum Target log
232710 04/17/08 13:39 R. Michaels BB wire chamber ok now.
232707 04/17/08 13:30 R. Michaels magnet screens
232704 04/17/08 13:07 R. Michaels NMR to probe B
232703 04/17/08 13:05 R. Michaels boring scaler checks
232702 04/17/08 13:01 R. Michaels going to production mode
232700 04/17/08 12:53 R. Michaels rebooted ROCs - now Hyd elast.
232695 04/17/08 12:28 V. Sulkosky Energy Lock back on Hall C
232694 04/17/08 12:25 R. Michaels energy lock to hall C
232688 04/17/08 12:10 R. Michaels BB wire chamber problem
232684 04/17/08 11:36 R. Michaels ask for energy lock on hall A
232681 04/17/08 11:26 V. Sulkosky Beam energy Drifting again
232676 04/17/08 10:55 R. Michaels magnet recovery
232675 04/17/08 10:23 R. Michaels controlled access to fix magnets
232674 04/17/08 10:21 R. Michaels adaql3 monitor/mouse ok
232672 04/17/08 10:09 R. Michaels lost magnets
232670 04/17/08 10:03 R. Michaels computer problems / beam back
232667 04/17/08 10:00 R. Michaels 3255_BAD file
232664 04/17/08 09:50 jpchen Deuterium loop (loop 2) warmed up, LH2 fan still 29 Hz
232661 04/17/08 09:41 jpchen Warming up deuterium loop and increasing LH2 fan speed to 36 Hz
232660 04/17/08 09:09 R. Michaels spot move / move to Kine G
232656 04/17/08 08:42 R. Michaels move to alum. tgt.
232649 04/17/08 08:02 C. Fernandez Owl Shift Summary
232640 04/17/08 07:15 C. Fernandez Run 3256 analysis
232631 04/17/08 06:34 C. Fenandez- Beam current snapshot
232622 04/17/08 05:57 C. Fernandez Beam not very well
232615 04/17/08 05:08 C. Fernandez Beam off. Many trips
232612 04/17/08 04:56 C. Fernandez Gas sensor
232607 04/17/08 04:42 Jose Udias target screens
232596 04/17/08 03:57 adaq Hall A Tools snapshot
232595 04/17/08 03:57 adaq Beam Position and Current snapshot
232594 04/17/08 03:57 adaq Beam Energy in A and C snapshot
232569 04/17/08 01:58 C. Fernandez BPM Snapshot
232568 04/17/08 01:58 C. Fernandez Hall A tools snapshot
232567 04/17/08 01:57 C. Fernandez Hall A current snapshot
232562 04/17/08 01:29 C. Fernandez In Run 3228 the target is loop1 LH2
232545 04/17/08 00:02 S. Sirca Swing shift summary
Wednesday
232543 04/16/08 23:57 jinhuang Temporarily Change Golden Rootfile to Run 3213
232540 04/16/08 23:45 S. Sirca Wrong calculation of p_proton
232533 04/16/08 23:10 Ole Hansen Notified Computer Center about adaqfs lockd
232524 04/16/08 22:20 jinhuang online replay updated
232517 04/16/08 21:48 Seonho Choi ROC5 problem resolved without rebooting
232505 04/16/08 20:45 V. Sulkosky T3 and T7 Comparison
232497 04/16/08 19:57 adaq Hall A Tools snapshot
232496 04/16/08 19:57 adaq Beam Position and Current snapshot
232495 04/16/08 19:57 adaq Beam Energy in A and C snapshot
232488 04/16/08 19:05 B Norum Target status
232479 04/16/08 18:23 S. Sirca BB HV Slot 14
232478 04/16/08 18:22 S. Sirca Beam energy stripchart
232477 04/16/08 18:21 S. Sirca BPM/BCM stripcharts
232476 04/16/08 18:19 S. Sirca Hall A General Tools
232473 04/16/08 18:11 brads Quads were cycled (late entry)
232470 04/16/08 17:52 brads adaqfs NFS lockd problem
232463 04/16/08 17:26 jinhuang update on Bigbitelib, rebuild scripts
232462 04/16/08 17:21 V. Sulkosky T3 and T7 Test
232457 04/16/08 17:02 Moffit Comparison of T3 only and T7 only
232451 04/16/08 16:50 jpchen cryotarget status last night and now
232447 04/16/08 16:33 Seonho Choi Negative dead time during run 3183
232442 04/16/08 16:02 LeRose Shift Summary
232432 04/16/08 15:27 Moffit Deatimes from T3 and T7 from run 3174
232422 04/16/08 14:46 Higinbotham Check of T7 Trigger vs T3 with Partial S1
232399 04/16/08 12:51 jpchen Adjusted LH2 to be 20K, 29Hz, lowered flow to LD2 loop.
232396 04/16/08 12:48 LeRose Beam Energy/ BB HV
232395 04/16/08 12:42 LeRose Hall A Tools BPM/BCM
232386 04/16/08 11:57 adaq Hall A Tools snapshot
232385 04/16/08 11:56 adaq Beam Position and Current snapshot
232384 04/16/08 11:56 adaq Beam Energy in A and C snapshot
232379 04/16/08 11:41 LeRose Hall A Energy Locks ON
232378 04/16/08 11:35 C. Fernandez Warning in ROC5
232371 04/16/08 11:01 C. Fernandez ROC 5 crach in Run 3153. All OK now
232359 04/16/08 10:26 V. Sulkosky Changed Golden Rootfile to Run 3139
232358 04/16/08 10:23 V. Sulkosky Late Entry: S1 1st and 6th Paddles On
232353 04/16/08 10:04 brads Chamber 2, U13 seems OK now
232344 04/16/08 09:14 M. Stevens/T Rt Q3 He lvl gauge false readback during ESR recovery
232337 04/16/08 08:44 LeRose Beam Energy Strip Chart
232334 04/16/08 08:37 LeRose Beam is back
232333 04/16/08 08:29 LeRose. Beam Off
232330 04/16/08 08:16 M.Shabestari Owl Shift Summary
232326 04/16/08 08:01 M.Shabestari Hall A Tools snapshot
232325 04/16/08 08:00 M.Shabestari BCM and BPM snapshots
232316 04/16/08 07:19 M.Shabestari HV on S4, S6, S7, and S14 tripped
232315 04/16/08 07:13 M.Shabestari Target moved to LH2
232312 04/16/08 06:57 M.Shabestari Target screens
232309 04/16/08 06:34 M.Shabestari Beam energy stripchart
232307 04/16/08 06:32 M.Shabestari Target moved to Al
232305 04/16/08 06:25 M.Shabestari Radiation monitor snapshot
232298 04/16/08 05:42 Moffit ROC9 crashes...
232295 04/16/08 05:28 Moffit Moved hvgui to adaql1
232292 04/16/08 05:15 Moffit Access activity
232291 04/16/08 05:06 M.Shabestari Requested controlled access
232285 04/16/08 04:32 M.Shabestari BigBite HV and currents
232284 04/16/08 03:56 adaq Hall A Tools snapshot
232283 04/16/08 03:56 adaq Beam Position and Current snapshot
232282 04/16/08 03:56 adaq Beam Energy in A and C snapshot
232277 04/16/08 03:40 khem roc9 reboot itself
232267 04/16/08 03:02 M.Shabestari BCM and BPM snapshots
232266 04/16/08 03:00 M.Shabestari Target moved to carbon
232257 04/16/08 02:23 M.Shabestari Target moved to Ta
232256 04/16/08 02:16 M.Shabestari Target---> Empty---> Holey carbon
232255 04/16/08 01:55 M.Shabestari Harp scan results
232254 04/16/08 01:19 M.Shabestari Hall A Tools snapshot
232253 04/16/08 00:57 M.Shabestari Target moved to BeO
232251 04/16/08 00:21 ycwang Target screens
232250 04/16/08 00:19 ycwang Target screens
232248 04/16/08 00:14 Yi Qiang Swing Shift Summary
Tuesday
232247 04/15/08 23:16 Xiaohui Reboot iocha5
232246 04/15/08 19:56 adaq Hall A Tools snapshot
232245 04/15/08 19:56 adaq Beam Position and Current snapshot
232244 04/15/08 19:56 adaq Beam Energy in A and C snapshot
232243 04/15/08 18:26 brads Chamber2 U13 follow-up
232242 04/15/08 18:16 jpchen cryotarget cool-down and conditions
232241 04/15/08 18:13 brads Double checked BB shielding coverage from target/beamline
232238 04/15/08 17:58 Yi Qiang HRS Magnets Ready
232237 04/15/08 17:13 Yi Qiang BigBite Detector HV OFF
232236 04/15/08 17:10 M. Stevens/T Rt Q3 He lvl gauge false readback during ESR recovery
232234 04/15/08 15:59 Saha Shift Summary
232232 04/15/08 15:42 jpchen ESR tripped at noon time. Taregt warmed up. Recovering now.
232231 04/15/08 15:02 Jose Udias MCC prevision for beam back
232230 04/15/08 14:32 R. Michaels D3 L-HRS fixed (I think)
232229 04/15/08 12:47 brads Topped off He Bags.
232228 04/15/08 12:23 Jose Udias ESR crash
232227 04/15/08 12:12 V. Sulkosky ESR Has Crashed
232226 04/15/08 11:56 adaq Hall A Tools snapshot
232225 04/15/08 11:56 adaq Beam Position and Current snapshot
232224 04/15/08 11:56 adaq Beam Energy in A and C snapshot
232223 04/15/08 11:48 Moffit T5 and T6 now use T7 instead of T3
232209 04/15/08 10:51 R. Michaels D3 on L-HRS double pulsing
232204 04/15/08 10:42 V. Sulkosky L-HRS Momentum and Angle
232202 04/15/08 10:09 spiegel/tech Spectrometer move
232201 04/15/08 09:26 R. Michaels u3 updated in scaler.map
232200 04/15/08 08:46 R. Michaels scaler checks
232192 04/15/08 08:04 C. Fernandez Owl Shift Summary
232161 04/15/08 03:56 adaq Hall A Tools snapshot
232160 04/15/08 03:56 adaq Beam Position and Current snapshot
232159 04/15/08 03:56 adaq Beam Energy in A and C snapshot
232148 04/15/08 02:42 B. Hahn Target Screens
232143 04/15/08 02:15 C. Fernandez BPM snapshot
232142 04/15/08 02:14 C. Fernandez Hall A tools
232141 04/15/08 02:04 C. Fernandez Beam current very unstable
232136 04/15/08 01:36 B.Hahn target move to LH2+coll
232127 04/15/08 00:41 C. Fernandez In Run 3069 Target is DUMMY
232126 04/15/08 00:40 C. Fernandez Target is DUMMY
232123 04/15/08 00:32 B.Hahn Run 3067
232117 04/15/08 00:14 B. Hahn moving to dummy target
232115 04/15/08 00:05 Kalyan Swing shift summary
Monday
232103 04/14/08 23:00 Kalyan BPM strip chart
232090 04/14/08 21:15 D. McNulty ROC5 died and was revived
232081 04/14/08 20:18 Kalyan Beam energy strip chart
232080 04/14/08 20:17 Kalyan Beam energy strip chart
232079 04/14/08 20:16 Kalyan HallA General Tools
232076 04/14/08 19:56 adaq Hall A Tools snapshot
232075 04/14/08 19:56 adaq Beam Position and Current snapshot
232074 04/14/08 19:56 adaq Beam Energy in A and C snapshot
232049 04/14/08 17:23 Kalyan BigBite HV and currents
232044 04/14/08 16:34 ycwang Target screens
232040 04/14/08 16:06 jpchen Day shift summary
232039 04/14/08 15:59 jpchen RF work, beam off
232038 04/14/08 15:56 jpchen back to beam permit
232037 04/14/08 15:32 jpchen requested controlled access to change gas
232032 04/14/08 15:12 jpchen One of the bubblers not bubbling.
232027 04/14/08 14:47 jpchen BigBite chamber HVs
232026 04/14/08 14:46 jpchen BPM/BCM
232025 04/14/08 14:46 jpchen Hall A tools
232021 04/14/08 14:37 Jose Udias Target screens
231994 04/14/08 11:56 adaq Hall A Tools snapshot
231993 04/14/08 11:56 adaq Beam Position and Current snapshot
231992 04/14/08 11:56 adaq Beam Energy in A and C snapshot
231973 04/14/08 10:07 jpchen Beam current stability
231965 04/14/08 09:26 Ole Hansen ROC5 reboot
231950 04/14/08 08:13 M.Shabestari Owl Shift Summary
231933 04/14/08 06:37 M.Shabestari BCM and BPM Snapshots
231924 04/14/08 05:42 M.Shabestari Beam Energy Stripchart
231923 04/14/08 05:41 M.Shabestari Radiation stripchart
231920 04/14/08 05:38 M.Shabestari Hall A Tools snapshot
231917 04/14/08 05:25 B.Hahn Target Screens
231904 04/14/08 03:56 adaq Hall A Tools snapshot
231903 04/14/08 03:56 adaq Beam Position and Current snapshot
231902 04/14/08 03:56 adaq Beam Energy in A and C snapshot
231879 04/14/08 01:32 M.Shabestari BigBite HV snapshot
231878 04/14/08 01:30 M.Shabestari Hall A Tools snapshot
231867 04/14/08 00:27 M.Shabestari BCM and BPM snapshots
231860 04/14/08 00:06 Kalyan swing shift summary
231858 04/14/08 00:03 Kalyan BigBite HV screens
Sunday
231857 04/13/08 23:57 Kalyan ROC5 rebooted - ADC errors
231842 04/13/08 22:29 Kalyan BPM strip chart
231829 04/13/08 21:07 Moffit Re: F1TDC warnings
231820 04/13/08 20:12 Kalyan ROC5 rebooted again
231818 04/13/08 20:11 Moffit Small change to hadron_ts.crl.
231816 04/13/08 19:58 Kalyan ROC 5 rebooted -- showing TDc errors
231814 04/13/08 19:55 adaq Hall A Tools snapshot
231813 04/13/08 19:55 adaq Beam Position and Current snapshot
231812 04/13/08 19:55 adaq Beam Energy in A and C snapshot
231810 04/13/08 19:44 Kalyan BigBite MWDC HV chart
231805 04/13/08 19:01 Kalyan Raster size -- spot++
231795 04/13/08 18:07 Kalyan BCM strip chart
231787 04/13/08 17:22 Kalyan Hall A General Tools screen
231783 04/13/08 17:03 Kalyan BPM chart
231777 04/13/08 16:13 Kalyan Target screens
231774 04/13/08 16:03 V. Sulkosky Day Shift Summary
231772 04/13/08 16:00 jose udias target screens (3/3)
231771 04/13/08 15:59 Jose Udias target screens (2)
231769 04/13/08 15:57 Jose Udias target screens
231765 04/13/08 15:35 V. Sulkosky Beam Instabilities and HV Trips
231760 04/13/08 14:57 V. Sulkosky BigBite HV GUI snapshot
231753 04/13/08 14:04 C. Fernandez ROC 5 stopped in run 2918
231751 04/13/08 13:55 adaq Beam Position and Current snapshot
231750 04/13/08 13:55 adaq Beam Energy in A and C snapshot
231743 04/13/08 12:55 V. Sulkosky Hall A Tools snapshot
231718 04/13/08 10:23 V. Sulkosky Re: CODA Rebooted
231715 04/13/08 10:17 V. Sulkosky Beam Trips
231712 04/13/08 10:00 C. Fernandez CODA Rebooted
231705 04/13/08 09:12 V. Sulkosky BPMA Y Position Drifted
231692 04/13/08 08:06 dhamilton Owl Shift Summary
231688 04/13/08 07:59 M.Shabestari Target ----> LH2
231687 04/13/08 07:45 M.Shabestari Target ---->Holey carbon
231686 04/13/08 07:38 dhamilton Beam restored to hall - beam position was off
231682 04/13/08 06:52 M.Shabestari Target screens
231681 04/13/08 05:49 dhamilton Beam off - electron gun problems
231675 04/13/08 05:00 dhamilton Production running at 2.5 microamps
231666 04/13/08 03:49 dhamilton HRSl set to 1.007 MeV/c
231661 04/13/08 03:23 M.Shabestari Target moved to LH2
231660 04/13/08 03:16 dhamilton HRSl set to 1158 MeV/c
231654 04/13/08 02:58 M.Shabestari Target moved to carbon
231652 04/13/08 02:42 dhamilton Beam alignment studies finished
231649 04/13/08 02:28 M.Shabestari Target moved to Ta
231648 04/13/08 02:13 M.Shabestari Target moved to empty
231647 04/13/08 00:44 M.Shabestari Target moved to BeO
231646 04/13/08 00:44 dhamilton Beam restored to hall
Saturday
231641 04/12/08 21:41 Ardashev Shift Summary
231640 04/12/08 20:00 adaq Hall A Tools snapshot
231639 04/12/08 20:00 adaq Beam Position and Current snapshot
231638 04/12/08 20:00 adaq Beam Energy in A and C snapshot
231636 04/12/08 18:42 R. Lindgren Shift Summary
231634 04/12/08 17:46 Ardashev target charts & info
231622 04/12/08 14:54 Lindgren BigBite Detector HV
231621 04/12/08 14:53 R. Lindgren Hall A Tools
231620 04/12/08 14:53 mercado Cryo Target Charts
231619 04/12/08 14:48 brads BigBite moved to 48 degrees (late entry)
231618 04/12/08 14:38 R. Lindgren Beam still down
231617 04/12/08 14:33 brads MWDC gas bottle status
231615 04/12/08 14:15 brads Restored 'goonlana' rootfile symlinks to point to adaql10 work disk
231614 04/12/08 13:49 R. Lindgren Beam Down
231613 04/12/08 13:48 R. Lindgren Beam Permit
231612 04/12/08 13:36 Moffit Access Activity. No T4 bad for runs 2796-2873
231611 04/12/08 13:35 brads Re: wirechamber database fixed (from 01:40:40 on April 12, 2008)
231610 04/12/08 13:25 brads Chamber 2 U13 fixed itself
231609 04/12/08 13:24 brads Did not find shorted signal wire on Chamber 2 U13 connector
231602 04/12/08 12:31 brads He concentration followup
231601 04/12/08 12:24 R. Lindgren T3 and T4 trigger check
231599 04/12/08 12:19 A. Camsonne T3 and T4 rates looks equals
231598 04/12/08 12:15 adaq Radiation and Ozone snapshot
231597 04/12/08 12:15 adaq Hall A Tools snapshot
231596 04/12/08 12:15 adaq Beam Position and Current snapshot
231595 04/12/08 12:12 Jack Left Arm CAMAC crates
231591 04/12/08 12:08 R. Lindgren Portmapper down
231588 04/12/08 12:06 Jack Check for afterpulsing in the BB PMTs
231587 04/12/08 12:04 A. Camsonne adaqfs portmapper issue
231585 04/12/08 11:46 Ole Hansen adaql10 / adaqfs problems
231584 04/12/08 11:40 R. Lindgren Control Access
231582 04/12/08 11:11 atrig Trigger-Download
231580 04/12/08 11:00 A. Camsonne Left S2 paddle seems OK now
231579 04/12/08 10:50 brads He concentration in/around flangle
231577 04/12/08 10:24 atrig Trigger-Download
231576 04/12/08 09:36 Moffit LHRS p0=1.194 GeV/c
231574 04/12/08 09:31 Moffit LHRS S2 PMT still low
231573 04/12/08 09:28 Moffit Target to Tantalum
231570 04/12/08 09:16 Moffit Going to controlled access
231564 04/12/08 08:42 R. Lindgren Day shift starting
231557 04/12/08 08:11 dhamilton Owl Shift Summary
231546 04/12/08 07:15 dhamilton BB HV slot 14 screenshot
231544 04/12/08 07:13 dhamilton Beam energy stripchart screenshot
231532 04/12/08 05:59 Ardashev target charts & info
231512 04/12/08 04:15 adaq Radiation and Ozone snapshot
231511 04/12/08 04:15 adaq Hall A Tools snapshot
231510 04/12/08 04:15 adaq Beam Position and Current snapshot
231504 04/12/08 03:53 Ardashev target charts & info
231488 04/12/08 02:34 dhamilton Beam is back - BB HV trip
231485 04/12/08 01:40 khem wirechamber database fixed
231482 04/12/08 01:37 dhamilton Beam off
231465 04/12/08 00:05 brads Hot wire in X1 plane in run 2813 (and more?)
231464 04/12/08 00:03 E.Chudakov Swing shift summary
Friday
231460 04/11/08 23:51 vladimir LH2 6 cm target snapshot
231459 04/11/08 23:44 brads /adaql10/work1 work disk is gone (previous online replay rootfiles offline!)
231456 04/11/08 23:39 E.Chudakov BPM charts
231455 04/11/08 23:37 E.Chudakov Tool, beam screens
231454 04/11/08 23:33 brads MWDC currents at -1500V
231453 04/11/08 23:32 E.Chudakov Moller results
231446 04/11/08 22:53 jinhuang MWDC database updated
231442 04/11/08 22:41 E.Chudakov BB DC HV -50V
231440 04/11/08 22:37 brads MWDC voltage and currents at -1550V
231439 04/11/08 22:36 brads/xin MWDC voltage increased +50V during runs 2810--2812 for efficiency check
231434 04/11/08 22:13 vladimir LH2 target snapshot
231431 04/11/08 22:00 E.Chudakov HV BB DC is up by 50V
231428 04/11/08 21:36 brads Should re-analyze files with bad t0
231421 04/11/08 20:44 vladimir target LH2
231420 04/11/08 20:36 Xin new t0 for drift time fix
231419 04/11/08 20:15 adaq Radiation and Ozone snapshot
231418 04/11/08 20:15 adaq Hall A Tools snapshot
231417 04/11/08 20:15 adaq Beam Position and Current snapshot
231416 04/11/08 18:57 vladimir empty target
231415 04/11/08 17:48 brads MWDCs time zeros wrong?
231414 04/11/08 17:41 E.Chudakov Start Moller measurement
231412 04/11/08 17:37 vladimir target status
231409 04/11/08 16:57 E.Chudakov Stopped data taking
231401 04/11/08 16:02 bogdanw day shift summary
231398 04/11/08 15:30 ge jin HRS left problem
231391 04/11/08 12:23 Moffit Re: Access. T7 changes at weldment
231390 04/11/08 12:15 adaq Radiation and Ozone snapshot
231389 04/11/08 12:15 adaq Hall A Tools snapshot
231388 04/11/08 12:15 adaq Beam Position and Current snapshot
231387 04/11/08 12:13 Moffit restarted HVgui. using new version
231383 04/11/08 11:52 R. Michaels EDTM off/on
231382 04/11/08 11:51 R. Michaels access: MPS, scaler gate, T7
231379 04/11/08 11:10 ge jin target screen
231374 04/11/08 10:50 ge jin target chart
231373 04/11/08 10:44 bogdanw BPMs
231371 04/11/08 10:15 JinGe BPM and energy chart
231370 04/11/08 10:13 JinGe HV HRS left
231369 04/11/08 10:11 JinGe HV screens HRS-R
231361 04/11/08 09:02 bogdanw Tools
231360 04/11/08 09:00 bogdanw BB MWDC currents
231350 04/11/08 08:10 dhamilton Owl Shift Summary
231340 04/11/08 07:00 dhamilton Beam energy screenshot
231338 04/11/08 06:59 dhamilton Beam energy screenshot
231337 04/11/08 06:58 dhamilton BPM screenshot
231336 04/11/08 06:57 dhamilton BB HV slot 14 screenshot
231331 04/11/08 06:28 dhamilton Beam restored to hall
231330 04/11/08 05:45 B. Hahn Target Screens
231328 04/11/08 04:03 adaq Radiation and Ozone snapshot
231327 04/11/08 04:03 adaq Hall A Tools snapshot
231326 04/11/08 04:00 adaq Beam Position and Current snapshot
231325 04/11/08 03:53 dhamilton Another update
231324 04/11/08 03:51 adaq no_key
231323 04/11/08 02:56 adaq Radiation and Ozone
231321 04/11/08 02:48 brads General Tools
231320 04/11/08 02:41 dhamilton Update
231319 04/11/08 02:37 adaq Hall A Tools
231317 04/11/08 02:11 dhamilton Beam down due to vacuum problems
231312 04/11/08 01:27 Moffit iocha5 reboot (BCM)
231299 04/11/08 00:04 fissum 2008-04-11 16:00 - 24:00 shift summary
Thursday
231288 04/10/08 23:08 Adaq snapshot -- beam-energy stripchart
231287 04/10/08 23:06 Adaq snapshot -- BPM and BCM stripchart
231256 04/10/08 19:14 adaq snapshot -- BB HV slot14
231255 04/10/08 19:06 Adaq snapshot -- Hall A General Tools
231248 04/10/08 18:29 brads How to extract photos from CODA
231240 04/10/08 17:26 jinhuang bigbite software lib updated
231224 04/10/08 15:53 R. Michaels white alarm on LeCroy HV
231221 04/10/08 15:37 gilman target for shift checklist
231218 04/10/08 15:17 R. Michaels production started
231216 04/10/08 15:04 R. Michaels cycling Q2,Q3 on L-HRS
231215 04/10/08 15:00 R. Michaels L-HRS dipole problem/solution
231212 04/10/08 14:10 gilman changed target
231210 04/10/08 14:02 R. Michaels L-HRS momentum change
231209 04/10/08 13:59 Norum Calibrations/normalizations
231208 04/10/08 13:56 R. Michaels more Ta elastic
231206 04/10/08 13:49 gilman target moved
231204 04/10/08 13:45 R. Michaels re: deadtime with T4
231203 04/10/08 13:37 R. Michaels deadtime rise
231202 04/10/08 13:34 R. Michaels adjust ps3,ps4 for elastic C12 run 2734
231199 04/10/08 13:26 R. Michaels re: Calibration run
231198 04/10/08 13:24 R. Michaels recent events
231190 04/10/08 13:06 gilman target change
231186 04/10/08 12:46 R. Michaels prescales for elastic Ta run 2728
231184 04/10/08 12:41 Norum Calibration run
231183 04/10/08 12:39 R.Lindgren BigBite angle slightly off
231182 04/10/08 12:39 gilman target moved
231179 04/10/08 12:24 Moffit Access activity
231178 04/10/08 12:23 spiegel/Tech Move LEFT Spectrometer
231177 04/10/08 12:01 gilman energy
231176 04/10/08 12:00 gilman energy change last evening
231173 04/10/08 11:22 R. Michaels Scaler checkout, scaler.map update
231172 04/10/08 10:57 gilman snapshot of bigbite power supply
231171 04/10/08 10:42 gilman setting up access
231159 04/10/08 09:26 R. Michaels reboot ROC5
231144 04/10/08 08:07 dhamilton E-dE plots (kinematics C)
231141 04/10/08 07:59 dhamilton Owl Shift Summary
231123 04/10/08 05:56 dhamilton Rise in deadtime - possible network problems
231092 04/10/08 03:07 B. Hahn Target Screens
231090 04/10/08 02:57 dhamilton New golden run for replay
231077 04/10/08 01:43 dhamilton BB HV 14 screenshot
231076 04/10/08 01:42 dhamilton Beam energy screenshot
231075 04/10/08 01:40 dhamilton BPM and BCM screenshots
Wednesday
231048 04/09/08 23:16 jinhuang script bug fix
231036 04/09/08 22:19 Xiaohui Target Position
231035 04/09/08 22:19 Xiaohui Target charts
231020 04/09/08 20:59 Norum Accelerator rf
231018 04/09/08 20:42 Higinbotham Hall A Access Today (late entry)
231017 04/09/08 19:54 Norum accelerator rf
231012 04/09/08 18:54 Norum HV trip
231005 04/09/08 18:16 doug Kinematics C Production
231000 04/09/08 17:55 brads He balloon camera now on Monitor and logged to CODA event 138
230999 04/09/08 17:55 brads LHRS angle camera logged to CODA event 135
230990 04/09/08 17:04 Moffit re: Updated coincidence window
230983 04/09/08 16:22 Higinbotham Day Shift Summary
230979 04/09/08 15:58 Moffit Another coincidence window change
230953 04/09/08 14:37 Higinbotham Quads Cycled
230947 04/09/08 13:02 Higinbotham End Kinematics B
230935 04/09/08 11:21 Moffit Re: coincidence window change
230929 04/09/08 10:47 Moffit Coincidence window change
230924 04/09/08 10:13 Higinbotham No BigBite HV Trips
230917 04/09/08 09:48 Higinbotham Beam Energy Plot After Trip
230914 04/09/08 09:34 Higinbotham Compton Position
230913 04/09/08 09:34 Higinbotham Compton Position
230912 04/09/08 09:32 Higinbotham BigBite MWDC HV
230911 04/09/08 09:30 Higinbotham BCM Striptool
230910 04/09/08 09:29 Higinbotham Compton Rates 5 uA
230901 04/09/08 08:39 M.Shabestari Owl Shift Summary
230898 04/09/08 08:36 Moffit Restarted HV gui.
230895 04/09/08 08:21 Moffit ROC5 rebooted. ADC errors
230891 04/09/08 07:56 M.Shabestari Target screens
230860 04/09/08 05:32 M.Shabestari Radiation stripchart
230851 04/09/08 05:06 M.Shaestari Target screens
230850 04/09/08 05:04 M.Shabestari Harebooted ROC5
230849 04/09/08 05:02 M.Shabestari Restarted hvgui
230841 04/09/08 04:26 M.Shabestari Changed set hreshold for dE
230836 04/09/08 04:02 M.Shabestari Hall A tools snapshot
230823 04/09/08 03:13 M.Shabestari Status
230819 04/09/08 02:55 M.Shabestari BPM and BCM snapshots
230817 04/09/08 02:50 M.Shabestari Target moved to LH2
230816 04/09/08 02:39 M.Shabestari T1 & T2 (target: Holey Carbon)
230815 04/09/08 02:28 M.Shabestari Target moved to Holey Carbon
230814 04/09/08 02:22 M.Shabestar T1 & T2 (target: empty)
230813 04/09/08 02:16 M.Shabestari Target moved to empty
230812 04/09/08 02:07 M.Shabestari Harp scan results
230810 04/09/08 01:53 M.Shabestari Target moved to BeO
230809 04/09/08 01:23 M.Shabestari Beam restored
230803 04/09/08 00:05 dhamilton Swing shift summary
Tuesday
230802 04/08/08 23:50 dhamilton Status - update
230801 04/08/08 21:41 E.Chudakov Moller status
230799 04/08/08 21:24 Moffit Run plan for Tuesday Swing/Wednesday Owl
230797 04/08/08 21:00 jinhuang Replay script now supports splitted CODA files
230796 04/08/08 16:48 Higinbotham Beam Energy from Ta Run 2535
230795 04/08/08 16:46 jinhuang Bigbite software update
230794 04/08/08 16:20 E.Chudakov Day shift summary
230791 04/08/08 15:56 Higinbotham Photo of LHRS Floor Mark
230790 04/08/08 14:06 E.Chudakov MCC update
230789 04/08/08 12:12 Higinbotham BigBite Chamber HV Set To 1500V
230788 04/08/08 11:41 E.Chudakov Beam off - vacuum problems
230787 04/08/08 10:33 E.Chudakov Start Moller measurement
230781 04/08/08 10:02 E.Chudakov Flipped MPS signal in the Moller DAQ
230761 04/08/08 08:18 M.Shabestari Owl Shift Summary
230759 04/08/08 08:14 M.Shabestari ROC5 error
230757 04/08/08 07:50 M.Shabestari Deadtime increased
230754 04/08/08 07:36 M.Shabestari Target screens
230753 04/08/08 07:32 M.Shabestari BCM & BPM snapshots
230747 04/08/08 07:16 M.Shabestari Beam trip ----> HV trip of S14
230744 04/08/08 06:57 M.Shabestari ROC5 error
230734 04/08/08 06:01 M.Shabestari HV and BCM snapshots
230727 04/08/08 05:21 M.Shabestari Hall A tools snapshot
230726 04/08/08 05:18 M.Shabestari HV trip: S14
230721 04/08/08 05:00 M.Shabestari Target screens
230703 04/08/08 03:31 M.Shabestari HV trip
230700 04/08/08 03:17 M.Shabestari ROC5 error
230685 04/08/08 02:05 M.Shabestari HV trip
230674 04/08/08 01:32 M.Shabestari HV trip
230672 04/08/08 01:27 M.Shabestari BPM and BCM snapshots
230670 04/08/08 01:20 M.Shabestari HV trip
230663 04/08/08 00:50 M.Shabestari HV trip
230662 04/08/08 00:33 M.Shabestari Target moved to LH2
230656 04/08/08 00:09 jinhuang All time stamps in db_run.dat are fixed
230653 04/08/08 00:01 dhamilton Swing shift summary
Monday
230650 04/07/08 23:57 A. Camsonne Left arm U2 plane seems fine after TDC change
230648 04/07/08 23:51 dhamilton HRSl set to 1194 MeV/c and 16.5 degrees
230647 04/07/08 23:29 dhamilton BB coil current back at 390A
230641 04/07/08 23:08 jinhuang Problem Fixed: time stamps in db_run.dat
230639 04/07/08 23:03 dhamilton BB coil current 200A
230632 04/07/08 22:40 A. Camsonne MWDC threshold set to 5V
230625 04/07/08 22:22 A. Camsonne threshold MWDC set to 10 V
230614 04/07/08 21:49 dhamilton BB MWDC HVGui at 0.5 microamps, 1450V, BB magnet off
230579 04/07/08 19:55 brads/doug Balloon work (late entry)
230566 04/07/08 19:10 dhamilton Beam restored to hall
230562 04/07/08 18:18 dhamilton Hall status - beam permit
230561 04/07/08 18:11 brads Balloon cam hooked up
230560 04/07/08 17:32 dhamilton Hall status changed to controlled access
230558 04/07/08 16:03 jpchen day shift summary
230557 04/07/08 16:02 jpchen day shift summary
230555 04/07/08 14:34 jpchen target screens
230554 04/07/08 14:23 Ole Hansen libBigBite recompiled with optimizations
230553 04/07/08 13:55 Ole Hansen Podd recompiled with optimizations
230552 04/07/08 13:51 Ole Hansen adaql5 is 1 core only
230551 04/07/08 12:53 jpchen Controlled access activities completed
230550 04/07/08 12:23 Moffit Added inverted MPS
230549 04/07/08 12:22 Moffit Fixed x2 rate in LHRS T1
230548 04/07/08 11:28 A. Camsonne TDC ROC3 slot 4 replaced
230544 04/07/08 10:08 A. Camsonne u2 wire inefficiency
230543 04/07/08 09:46 jpchen L_HRS set for kinematics B
230542 04/07/08 09:34 jpchen change to kinematics B
230541 04/07/08 09:15 jpchen Beam off, wait for controlled access
230540 04/07/08 09:11 Xiaohui Target: empty
230539 04/07/08 08:58 jpchen BigBite HV trip
230535 04/07/08 08:49 jpchen BCM OLO2 spike
230532 04/07/08 08:34 M.Shabestari Owl Shift Summary
230531 04/07/08 08:31 jpchen HV trip
230529 04/07/08 08:20 khem big hv trip, all setting gone
230527 04/07/08 08:01 R. Lindgren Helium Flow
230526 04/07/08 07:57 M.Shabestari HV trip: S1, S4, S14
230522 04/07/08 07:42 M.Shabestari Htr_aft @ LOOP2
230520 04/07/08 07:15 M.Shabestari High Htr_aft in LOOP2
230519 04/07/08 07:08 M.Shabestari HV trip: S1, S5, S14
230517 04/07/08 06:53 M.Shabestari BPM and BCM snapshots
230512 04/07/08 06:04 M.Shabestari HV trip again
230510 04/07/08 05:30 M.Shabestari HV trip
230504 04/07/08 04:55 M.Shabestari BPM and BCM snapshots
230501 04/07/08 04:14 M.Shabestari Target Screens
230500 04/07/08 04:13 M.Shabestari Hall A tools & BCM snapshots
230499 04/07/08 04:12 M.Shabestari HV snapshot
230498 04/07/08 04:10 M.Shabestari HVs Trip
230494 04/07/08 03:34 M.Shabestari HV trip
230491 04/07/08 02:52 M.Shabestari HV Trip
230487 04/07/08 02:45 Moffit x2 difference in rate in T1s!!
230484 04/07/08 02:38 M.Shabestari HV trip
230483 04/07/08 02:33 Moffit Left Arm HV snapshot
230481 04/07/08 02:27 brads He Ballon status (getting low)
230475 04/07/08 01:53 M.Shabestari Tools and BPM screenshots
230474 04/07/08 01:34 brads Rear chamber problem summary
230469 04/07/08 01:05 A. Camsonne BB MWDC chamber HV
230468 04/07/08 00:58 Moffit Updated mwdc database
230467 04/07/08 00:46 A. Camsonne Plan for tonight
230465 04/07/08 00:43 A. Camsonne Left HRS NMR dipole could not lock on probe A
230463 04/07/08 00:40 Moffit Access activity
230462 04/07/08 00:37 dhamilton HRSl set back to 960 MeV/c
230458 04/07/08 00:05 dhamilton Swing Shift Summary
Sunday
230455 04/06/08 23:49 dhamilton BB coil current set to 0A
230453 04/06/08 23:45 dhamilton HRSr P0 regulation for Q2 on
230450 04/06/08 23:39 dhamilton BCM and Radiation screenshot
230449 04/06/08 23:36 dhamilton Tools and BPM screenshot
230448 04/06/08 23:33 brads He and Ethane bottle status
230438 04/06/08 22:50 dhamilton HRSl set to 1194 MeV/c
230436 04/06/08 21:23 dhamilton Hall Status - controlled access
230433 04/06/08 21:02 brads Automatic logging of LHRS angle
230426 04/06/08 18:50 Lindgren Day shift summary
230423 04/06/08 18:41 Moffit CODA file splitting enabled.
230422 04/06/08 18:30 jinhuang update of mwdc database and replay script
230415 04/06/08 17:42 fissum BigBite Trigger supervisor Deadtime 100%
230411 04/06/08 17:15 Higinbotham Photo of HRS Angle @ 20.5
230402 04/06/08 16:19 R. Lindgren Shift Summary
230401 04/06/08 16:19 R.Lindgren BigBite singles
230391 04/06/08 15:30 R. Lindgren Status
230380 04/06/08 14:37 R. Michaels MPS signal checked
230363 04/06/08 12:37 R. Lindgren pion rejector
230350 04/06/08 11:27 Higinbotham T4 Trigger Comment
230347 04/06/08 11:09 R. Lindgren T4
230345 04/06/08 11:08 R. Lindgren S1 intentionally turned of
230343 04/06/08 11:05 R. LIndgren Status of Running
230330 04/06/08 09:43 R. Michaels re: helicity
230324 04/06/08 09:27 Higinbotham Plots with S1 outer paddes off
230322 04/06/08 09:23 Higinbotham Reduced S1 Acceptance
230321 04/06/08 09:18 Higinbotham Helicity Problem?
230307 04/06/08 07:38 Higinbotham New Coinc Window Data
230306 04/06/08 07:37 M.Shabestari Target screens
230299 04/06/08 06:40 Moffit Coincidence window change
230297 04/06/08 06:37 Higinbotham S1 Outer Paddles Off
230283 04/06/08 04:30 Moffit Removed unused HV crate from GUI
230276 04/06/08 04:03 M.Shabestari Target screens
230266 04/06/08 03:16 Moffit ROC5 out of memory? Rebooted.
230243 04/06/08 01:30 brads THaNormAna deadtime calculations available
230232 04/06/08 00:45 M.Shabestari Target screens
230231 04/06/08 00:39 Moffit pedsup for ROC3 and ROC4 didn't help, datafile NOT splitting
230221 04/06/08 00:03 dhamilton Swing Shift Summary
230220 04/06/08 00:03 brads To all analysis code people -- we have no quad-core machines
230219 04/06/08 00:02 khem wire chamber 2u'11
Saturday
230217 04/05/08 23:56 Moffit Turned on pedestal suppression for ROC3 and ROC4
230216 04/05/08 23:49 dhamilton BB HV GUI crashed
230205 04/05/08 22:58 jinhuang Correction: No BigBite Sieve since run # 2353
230204 04/05/08 22:52 dhamilton BB HV trip
230203 04/05/08 22:51 dhamilton Right arm Q2 P0 regulation switched off for now
230199 04/05/08 22:41 dhamilton Right Arm Q2 power supply trip
230197 04/05/08 22:37 Moffit DB CVS created... used
230195 04/05/08 22:28 dhamilton HRS Right Q2 current at 0 A
230187 04/05/08 21:23 dhamilton Move to production kinematics
230186 04/05/08 21:19 dhamilton HRSl pcent set to 1104.9 MeV/c
230179 04/05/08 21:09 jinhuang MWDC crate map updated
230164 04/05/08 20:31 brads MWDC gas bottle pressure correction
230163 04/05/08 20:27 Moffit Controlled Access activity...
230162 04/05/08 20:25 jinhuang target->C
230155 04/05/08 19:54 dhamilton Raster off - 5 microamps on target
230154 04/05/08 19:43 dhamilton HRSl momentum set to 1194 MeV/c
230153 04/05/08 19:40 jinhuang target moved to Ta
230147 04/05/08 19:30 dhamilton Hall status - beam permit
230140 04/05/08 19:18 A. Camsonne Gas bottles status
230139 04/05/08 19:17 A. Camsonne Big bite sieve slit out
230132 04/05/08 18:48 dhamilton Move HRSl to 20.5 deg, pcent=960 MeV/c, BB coil I=390A
230129 04/05/08 18:37 Lindgren Shift Summary
230118 04/05/08 16:49 Moffit Increased event buffer size on ROC5
230117 04/05/08 16:44 jinhuang target ->LH2
230105 04/05/08 16:11 Moffit MWDC HV trip.
230092 04/05/08 16:01 Moffit FPP HV Off
230089 04/05/08 15:46 R. Lindgren LH2 Runs
230088 04/05/08 15:45 Moffit ER1 crashed... restarted CODA
230082 04/05/08 14:58 brads TECH: RHRS helium flows are unstable.
230081 04/05/08 14:51 brads Restarted BB HV gui
230080 04/05/08 14:41 brads High load on adaql1 -- killed unused etserver processes
230079 04/05/08 14:40 R. Michaels etserver killed off
230072 04/05/08 14:20 brads Raster -> 1.5x1 (MCC units) for production
230071 04/05/08 14:10 brads Beam position and raster is excellent.
230064 04/05/08 14:01 brads Increased raster to 4.5x3 (MCC units) to check position on holey carbon
230059 04/05/08 13:53 Ge Jin carbon w/ hole
230058 04/05/08 13:43 R. Lindgren Left HRS
230057 04/05/08 13:42 brads Target -> Empty for initial beam delivery
230056 04/05/08 13:23 R. Michaels datamon mods
230053 04/05/08 13:05 R. Michaels helicity scalers fixed (L-HRS)
230052 04/05/08 13:02 R. Lindgren No beam.
230049 04/05/08 12:49 brads Back to beam permit (MCC still working on other Halls)
230048 04/05/08 12:46 brads Suggested addition to end-of-run replay output
230043 04/05/08 12:33 brads Adjusted MWDC flow rate
230042 04/05/08 12:33 brads Opportunistic controlled access
230041 04/05/08 11:43 brads/lindgr Issue with inserting BB sieve slit
230040 04/05/08 11:39 brads MWDC gas flow can be turned down
230039 04/05/08 11:34 brads Camera server on hacweb1 back online
230038 04/05/08 11:20 R. Lindgren Control access
230037 04/05/08 10:56 W. Luo Load HV from file hv_70025.new, corrected the hot rows
230036 04/05/08 10:03 R. Michaels re: datamon
230035 04/05/08 09:52 R. Lindgren Roc 5 Crash and Hall A acces
230027 04/05/08 09:17 Higinbotham Owl Shift Summary
230023 04/05/08 08:46 Higinbotham Datamon Deadtime Test
229999 04/05/08 07:41 Higinbotham Left Arm to 20.5 Degrees
229995 04/05/08 07:30 vladimir set Carbon Target
229994 04/05/08 07:28 Higinbotham dE Trigger Threshold Returned
229980 04/05/08 06:04 A. Camsonne MWDC HV trip
229979 04/05/08 05:47 vladimir LD2 target screen
229977 04/05/08 05:45 vladimir LD2 target screen
229975 04/05/08 05:13 A. Camsonne Odd value of dead time
229974 04/05/08 04:55 vladimir JT loop 2 increased
229971 04/05/08 04:50 A. Camsonne High deadtime value
229969 04/05/08 04:48 A. Camsonne dE threshold trigger set to 560
229967 04/05/08 04:43 vladimir cryo target screen
229966 04/05/08 04:39 A. Camsonne Problem setting threshold
229964 04/05/08 04:30 vladimir LD2 screen
229962 04/05/08 04:20 A. Camsonne ROC5 in error condition
229961 04/05/08 04:18 vladimir LD2 target status
229958 04/05/08 04:07 A. Camsonne ROC 5 warning
229957 04/05/08 03:35 A. Camsonne Missing TDC run 2309
229953 04/05/08 03:12 A. Camsonne ROC5 rebooted by itself
229950 04/05/08 02:03 A. Camsonne Run 2308 HV off on one DE channel
229947 04/05/08 01:46 A. Camsonne Scaler rates LD2 BB at 390 A
229946 04/05/08 01:45 A. Camsonne Bigbite current set to 0
229942 04/05/08 01:38 vladimir LD2 6 cm target
229941 04/05/08 01:31 A. Camsonne Bigbite rates with magnet off
229940 04/05/08 01:25 A. Camsonne Bigbite current set to 0
229938 04/05/08 01:15 A. Camsonne Scaler rates LH2 2.5 uA
229937 04/05/08 01:07 A. Camsonne BB MWDC HV and current
229936 04/05/08 01:01 vladimir LH2 target status
229932 04/05/08 00:38 A. Camsonne BB Scintillators HV trip end of run 2304
229931 04/05/08 00:24 A. Camsonne HRS VDC efficiency
229929 04/05/08 00:06 E.Chudakov Shift summary
Friday
229927 04/04/08 23:48 jinhuang ROC5 rebooted
229926 04/04/08 23:45 xqian HV and threshold scan for LH2 at 3uA; background rate
229924 04/04/08 23:39 jinhuang ROC 5 rebooted
229923 04/04/08 23:24 jinhuang ROC5 reboot
229921 04/04/08 22:44 R. Michaels last scaler reading ok now
229920 04/04/08 22:42 jinhuang Reboot ROC 5 from hareboot26
229917 04/04/08 22:27 R. Michaels scaler problems / prob. solns.
229915 04/04/08 22:23 jinhuang ROC5 reboot
229914 04/04/08 22:06 E.Chudakov Radiation stripchart, Compton OTR
229912 04/04/08 22:05 E.Chudakov Tools, BB screen
229911 04/04/08 22:02 E.Chudakov Beam stripchart
229910 04/04/08 21:57 E.Chudakov LHRS sieve is out
229909 04/04/08 21:47 E.Chudakov BB magnet is back to 390A
229908 04/04/08 21:46 E.Chudakov Running with BB magnet off - failed
229907 04/04/08 20:25 khem online gui for bigbite now change from raw adc to adc with pedestal subtraction
229906 04/04/08 20:14 jinhuang Update on Replay Script
229905 04/04/08 19:05 Moffit inaccurate number for LHRS angle
229904 04/04/08 18:41 Moffit recompiled TS12 crl, re-downloaded
229903 04/04/08 18:22 E.Chudakov BB magnet is set to zero
229900 04/04/08 18:02 E.Chudakov BB Dc HV increased to 1550V
229897 04/04/08 17:50 jinhuang ROC 5 Reboot
229892 04/04/08 16:22 xqian Background on Carbon 5uA for different HV and threshold on chamber!!
229890 04/04/08 16:06 bogdanw shift summary
229889 04/04/08 15:44 bogdanw BB HV
229888 04/04/08 15:43 V. Sulkosky Added BigBite Gaussmeter to Loggers
229887 04/04/08 15:43 Adaq tools
229883 04/04/08 15:17 A. Camsonne MWDC HV on BB increased to 1500 V
229881 04/04/08 15:15 A. Camsonne Event sync at EB
229869 04/04/08 13:57 LeRose Target: 6 cm H2
229868 04/04/08 12:35 LeRose Empty Target
229865 04/04/08 12:08 LeRose Target Charts
229864 04/04/08 12:03 LeRose Target Screens
229860 04/04/08 09:17 bogdanw Right arm NMR
229859 04/04/08 08:35 bogdanw Hall status
229858 04/04/08 08:34 LeRose H2 target
229857 04/04/08 08:17 dhamilton Owl Shift Summary
229855 04/04/08 07:50 dhamilton BB MWDC HV switched off
229854 04/04/08 07:41 vladimir empty target
229853 04/04/08 07:34 dhamilton Position problem being investigated - moving to empty target
229852 04/04/08 07:21 dhamilton BB HV trip - crew chief contacted
229850 04/04/08 07:01 vladimir LH2 target
229849 04/04/08 06:54 dhamilton Spot position on BeO
229848 04/04/08 06:46 dhamilton Harp scan complete
229847 04/04/08 06:46 vladimir BeO target
229846 04/04/08 06:27 vladimirpty empty target
229845 04/04/08 06:25 vladimir empty target
229844 04/04/08 06:11 dhamilton Beam position problems - empty target for harp scan
229843 04/04/08 05:57 dhamilton BB HV trip - again
229841 04/04/08 05:43 dhamilton BB HV MWDC trip
229840 04/04/08 05:34 vladimir LH2 target status
229838 04/04/08 05:32 vladimir set LH2 6 cm target
229837 04/04/08 05:24 dhamilton Beam alignment
229836 04/04/08 05:02 vladimir set BeO target
229835 04/04/08 04:57 dhamilton Beam restored to hall - BB HV trip
229834 04/04/08 03:14 khem add pedestal(from run 2256) for bbite pmt to database
229833 04/04/08 02:19 dhamilton Beam problems persist
229832 04/04/08 01:33 dhamilton Elastic protons in BB MWDC (390A)
229831 04/04/08 01:18 dhamilton Beam down for a while
229828 04/04/08 00:31 dhamilton Move to elastic optics running
229825 04/04/08 00:10 jpchen Swing shift summary
229824 04/04/08 00:09 jpchen threshold to 12.5V
229820 04/04/08 00:02 jpchen threshold to 10V
Thursday
229818 04/03/08 23:56 jpchen threshold from 5 to 7.5 V
229815 04/03/08 23:49 jpchen beam back, resume threshold/HV scan
229813 04/03/08 22:26 jpchen Beam off
229812 04/03/08 22:24 jpchen BigBite magnet setting
229811 04/03/08 22:23 jpchen changed BigBite chamber HVs to 1450 V
229809 04/03/08 22:20 jpchen HRS HVs
229808 04/03/08 22:19 jpchen BigBite HVs
229807 04/03/08 22:18 jinhuang target screen
229806 04/03/08 22:18 jpchen BPM/BCM and Compton screens
229805 04/03/08 22:17 jpchen Tools screen
229804 04/03/08 22:16 jpchen Threshold 7.5 --> 5V
229800 04/03/08 22:09 jpchen threshold 10V -> 7.5 V.
229798 04/03/08 22:04 jpchen lowered threshold from 12.5 to 10 V with HV=1350V.
229795 04/03/08 21:30 Moffit New DB directory...
0 04/03/08 21:12 no_name no_key
229791 04/03/08 20:51 jinhuang Correction: raster is off since run 2264
229790 04/03/08 20:38 jpchen Threshold scan, beam status
229781 04/03/08 18:26 A. Camsonne Compton restarted
229778 04/03/08 18:04 jpchen threshold raised to 7.5V
229776 04/03/08 17:54 A. Camsonne Tracks in bigbite ? HV=1400 V and threshold 5V
229768 04/03/08 17:19 jpchen started LH2 elastic
229766 04/03/08 17:05 jinhuang moving target to LH2
229765 04/03/08 17:01 jpchen harp scan results
229764 04/03/08 16:57 jinhuang change to BeO target
229763 04/03/08 16:53 jpchen Beam back, harp done, to BeO now
229762 04/03/08 16:47 jinhuang Empty Target, 5uA
229760 04/03/08 15:58 Saha Shift Summary
229759 04/03/08 14:43 spiegel Big Bite Ballon change out
229758 04/03/08 14:13 V. Sulkosky ADC Helicity Added to Replay Files
229755 04/03/08 13:37 A. Camsonne VDC looks ok on cosmics
229753 04/03/08 12:50 brads Front chamber cathode HV connector loose at top -- fixed
229752 04/03/08 12:42 gilman status of getting beam back
229749 04/03/08 12:31 gilman empty target
229748 04/03/08 12:22 Jack BigBite gauss probe readout in EPICS
229743 04/03/08 11:59 Higinbotham 4He Bags and WC Gas Checked
229742 04/03/08 11:58 Higinbotham Sieve Clamps Removed
229741 04/03/08 11:56 Higinbotham Right HRS -> 105 Degrees
229736 04/03/08 09:32 M. Stevens TECH/Target chamber lamp replaced
229735 04/03/08 09:29 R. Michaels TWOARM_FB scripts fixed
229734 04/03/08 09:27 R. Michaels online deadtime calc. updated
229733 04/03/08 09:26 xqian No signal on 1-U2 and 1-X1 plane for chamber-1
229727 04/03/08 09:02 gilman target status
229725 04/03/08 08:06 dhamilton Owl Shift Summary
229722 04/03/08 07:51 vladimir target LH2 6 cm
229721 04/03/08 07:45 dhamilton BB MWDC calibration runs complete
229715 04/03/08 07:25 vladimir cryo target status
229712 04/03/08 07:15 dhamilton MWDC HV set to 1550V
229703 04/03/08 06:55 dhamilton MWDC HV set to 1500V
229697 04/03/08 06:45 dhamiltonq Beam restored to Hall
229696 04/03/08 06:37 dhamilton BB E and dE thresholds
229695 04/03/08 06:23 dhamilton Accelerator problems
229689 04/03/08 05:46 dhamilton MWDC HV set to 1450V
229681 04/03/08 05:24 dhamilton BB MWDC HV set to 1400V
229672 04/03/08 04:41 dhamilton Restart BB MWDC HV/threshold scan
229671 04/03/08 04:40 Moffit Controlled Access activity... Missing data from TDC
229670 04/03/08 04:32 dhamilton Hall status changed to Beam Permit
229665 04/03/08 04:00 dhamilton Hall at controlled access
229658 04/03/08 02:26 dhamilton ROC5 problem
229655 04/03/08 02:10 dhamilton Problems with ROC5 persist
229648 04/03/08 01:32 dhamilton Problem with ROC5
229643 04/03/08 01:22 R. Michaels re: end-of-run scalers
229639 04/03/08 01:00 vladimir Carbon Target
229637 04/03/08 00:53 dhamilton Start of BB MWDC scan
229634 04/03/08 00:12 jinhuang Loop1 JT valve closed by 3%
229633 04/03/08 00:08 jinhuang Loop1 alarm again, JT valve opened for another 2%
229632 04/03/08 00:08 V. Sulkosky Swing Shift Summary
229630 04/03/08 00:01 jinhuang Loop1 alarm, JT opened by 2%
Wednesday
229627 04/02/08 23:41 jinhuang mwdc database updated
229626 04/02/08 23:27 jinhuang move to LH2 target
229625 04/02/08 23:27 V. Sulkosky Raster Scan Holely Carbon
229622 04/02/08 23:19 V. Sulkosky Raster Scan holely Carbon
229612 04/02/08 22:54 Moffit Something wrong with the scaler information in the end of runs.
229611 04/02/08 22:50 Moffit Quick controlled access
229605 04/02/08 22:29 V. Sulkosky Hall in Beam Permit
229604 04/02/08 22:28 jinhuang Loop2 JT valve close by .5%
229603 04/02/08 22:15 V. Sulkosky Controlled Access
229602 04/02/08 22:00 jinhuang Loop2 JT valve close by .5%
229601 04/02/08 21:46 jinhuang Light in target chamber is down
229600 04/02/08 21:41 jinhuang target camera seems
229593 04/02/08 21:24 jinhuang Change JT Valve Settings
229590 04/02/08 21:16 V. Sulkosky Harp Scan
229589 04/02/08 21:12 jinhuang Move target to carbon with hole
229584 04/02/08 20:31 jinhuang Loop 2 JT valve closed by .5%
229583 04/02/08 20:28 jinhuang bigbitelib update
229582 04/02/08 20:17 V. Sulkosky MWDC HVs
229579 04/02/08 20:11 brads Balloon 'excursion' and MWDC trip info
229577 04/02/08 19:47 Moffit Access Activity - VME to Fastbus switch
229576 04/02/08 19:33 jinhuang Loop JT valve be closed by .5%
229575 04/02/08 19:06 jinhuang JT valves closed by .5%
229568 04/02/08 18:49 V. Sulkosky Helicity Database Added to E04-007 Podd Database Directory
229567 04/02/08 18:30 jinhuang Begin reducing heter power
229566 04/02/08 18:17 Kalyan BigBite Fastbus crates synchronization-che
229563 04/02/08 18:09 Moffit Changing thresholds for MWDC
229558 04/02/08 17:44 V. Sulkosky MWDC HV Off
229557 04/02/08 17:36 Moffit Added signals to database.
229556 04/02/08 17:31 V. Sulkosky BigBite HV Tripped Off
229555 04/02/08 16:45 Moffit Signals moved from ROC5
229550 04/02/08 16:25 E.Chudakov Moller polarimeter status
229544 04/02/08 15:34 Higinbotham Deuteron @ 23.5 Oscilation
229540 04/02/08 14:38 Ole Hansen G0 helicity class updated
229538 04/02/08 13:43 Higinbotham Target Motion GUI Fixed
229537 04/02/08 13:42 Higinbotham Loop2 Temp @ 23.0
229536 04/02/08 13:12 A. Camsonne Beam position stripchart
229535 04/02/08 13:09 A. Camsonne Compton stripchart
229534 04/02/08 13:05 A. Camsonne Beam radiation strip chart
229532 04/02/08 12:22 jpchen Loop 2: changed PID to temperature sensor HX_crnx
229531 04/02/08 11:58 Jack Reset RQ2
229530 04/02/08 11:45 R. Michaels re: deadtime
229529 04/02/08 09:56 Moffit Coincidence window looks good.
229528 04/02/08 09:20 A. Camsonne Access time for switch from VME to fastbus will start around 10AM
229524 04/02/08 08:11 dhamilton Owl Shift Summary
229521 04/02/08 07:44 M.Shabestari Target screens
229519 04/02/08 07:38 khem roc 6 boot
229512 04/02/08 07:01 khem run#2163(which has runcontrol problem)
229511 04/02/08 06:54 dhamilton Problem with ROC5 again - beam position unstable.
229510 04/02/08 06:52 khem message roc5 before reboot itself
229505 04/02/08 06:17 dhamilton BB HV and ROC5 problems
229504 04/02/08 06:12 dhamilton Start of elastic coincidence running
229502 04/02/08 06:03 M.Shabestari Target moved to H2
229500 04/02/08 05:55 khem runcontrol problem- roc snap shot
229499 04/02/08 05:49 khem runcontrol problem
229498 04/02/08 05:49 dhamilton BB MWDC HV turned off
229495 04/02/08 05:43 dhamilton BB E/dE gain and discriminator threshold calibration with QE C data
229493 04/02/08 05:33 dhamilton BB MWDC HV set to 1350V
229490 04/02/08 05:20 dhamilton BB MWDC HV set to 1300V
229472 04/02/08 02:35 M.Shabestari Loop2 HP Heater set to PID mode
229465 04/02/08 02:12 khem deadtime (scaler)
229464 04/02/08 02:11 khem deadtime (prescale.dat)
229463 04/02/08 02:10 khem deadtime 98%
229460 04/02/08 01:31 dhamilton Sieve slit runs finished, moving to two-arm running
229458 04/02/08 01:29 M.Shabestari Target: Carbon
229457 04/02/08 01:21 Moffit Attempted to increase Coin window
229455 04/02/08 00:48 Moffit Coincidence window looks narrow.
229453 04/02/08 00:41 Ole Hansen Bug re cuts on histograms fixed
229451 04/02/08 00:23 M.Shabestari Loop 2 HP Heater set to manual mode
229449 04/02/08 00:00 Annand Swing shift summary
Tuesday
229448 04/01/08 23:52 jinhuang modifying MWDC database
229446 04/01/08 23:44 Annand BB HV OFF
229445 04/01/08 23:38 Annand Insert Sieve slit LHRS
229444 04/01/08 23:35 Annand Insert Sieve slit LHRS
229443 04/01/08 23:30 Nilanga MWDC currents
229439 04/01/08 23:25 A. Camsonne H2 6cm 5uA
229437 04/01/08 23:24 Nilanga MWDC currents
229433 04/01/08 23:19 Nilanga MWDC current
229428 04/01/08 23:13 jinhuang Move to LH2 target
229426 04/01/08 23:09 Annand Moved to LH2 target
229425 04/01/08 22:57 Nilanga MWDC currents
229415 04/01/08 22:29 Nilanga MWDC currents
229414 04/01/08 22:26 A. Camsonne Warning message F1 ROC5
229412 04/01/08 22:25 Nilanga MWDC currents
229409 04/01/08 22:17 jinhuang target moved to carbon
229408 04/01/08 22:16 jinhuang target loop2 sensor recover again
229407 04/01/08 22:11 Annand Move non-holey carbon target
229406 04/01/08 22:03 jinhuang target loop2 manual heater power no respond
229405 04/01/08 21:56 Annand Rast scan holey target raster 1.5 width, 1.0 height
229401 04/01/08 21:48 Annand Raster scan holey carbon...just about right
229399 04/01/08 21:41 Annand Raster scan BPMBY from -1.7 to -2.1
229396 04/01/08 21:32 A. Camsonne Restarted CODA... ROC5 rebooted by itself
229394 04/01/08 21:20 Annand Holey target raster
229389 04/01/08 21:13 Annand Raster scan BeO
229388 04/01/08 21:13 A. Camsonne Raster setting
229387 04/01/08 21:12 jinhuang Move target to Carbon with hole
229386 04/01/08 21:09 Annand Holey Carbon Target
229378 04/01/08 20:55 Annand BeO Beam position Check
229377 04/01/08 20:52 jinhuang Loop2 PID
229376 04/01/08 20:07 jinhuang Target moved to BeO
229375 04/01/08 20:01 A. Camsonne BigBite makes no visible beam deflection on the dump at 650 A
229374 04/01/08 19:53 A. Camsonne BigBite current to 300 A
229373 04/01/08 19:52 A. Camsonne BigBite current to 200 A no change
229372 04/01/08 19:50 A. Camsonne BigBite current 100 A no change
229371 04/01/08 19:50 A. Camsonne Current to 30 uA to view the beam on the dump
229370 04/01/08 19:47 A. Camsonne BigBite power supply would not turn on
229367 04/01/08 18:40 Annand Harp Scan Completed
229365 04/01/08 17:24 jinhuang Physics Replay Update
229364 04/01/08 17:01 brads Added copper tape to MWDC ground connectors
229363 04/01/08 16:46 brads v767 issues
229361 04/01/08 16:19 jinhuang target status: loop2 sensor recovered
229358 04/01/08 15:59 jpchen Loop 2 PID and temperature sensor.
229357 04/01/08 15:43 jpchen Coolant flow vs fan frequency
229356 04/01/08 15:27 Annand Evidence for protons in BigBite
229354 04/01/08 14:59 Moffit Re: dE-9R rate
229348 04/01/08 14:37 Jack Helium bags for BigBite
229347 04/01/08 14:28 Jack Left Arm VDC cabling OK at Fastbus crate
229343 04/01/08 13:32 jpchen target status: loop 2 temp sensor problems
229336 04/01/08 12:00 V. Sulkosky Added Sieve Slit Variables to Left and Right HRS Replays
229334 04/01/08 11:55 jpchen target position control corrected
229333 04/01/08 11:41 jinhuang class BBDecData is added into bigbitelib
229322 04/01/08 09:24 xqian Dead time 10%
229318 04/01/08 08:49 xqian ROC5 reboot itself
229317 04/01/08 08:42 Lindgren Shift Summary
229309 04/01/08 07:36 Lindgren Summary BPM/BCM Rad Mon
229308 04/01/08 07:31 Lindgren shift summary tools and current
229306 04/01/08 07:10 A. Camsonne Power cycled ROC5
229304 04/01/08 06:59 khem roc5 warning message
229303 04/01/08 06:58 Lindgren Reduce raster size and turn on wire chambers
229301 04/01/08 06:55 A. Camsonne Test BB wirechamber C 1 uA
229297 04/01/08 06:43 M.Shabestari Target moved to carbon
229296 04/01/08 06:28 A. Camsonne Left VDC issue
229295 04/01/08 06:26 A. Camsonne Comparison run 2101 and 2102
229293 04/01/08 06:23 M.Shabestari Target moved to BeO
229291 04/01/08 05:40 Lindgren Partial Shift Summary
229286 04/01/08 04:40 M.Shabestari Target moved to Carbon
229285 04/01/08 04:24 M.Shabestari Target moved to LH2
229278 04/01/08 03:58 M.Shabestari Target moved to BeO
229275 04/01/08 03:50 khem rates run number 2098
229274 04/01/08 03:46 khem rates run# 2098 no target
229273 04/01/08 03:44 khem scaler run# 2098
229272 04/01/08 03:42 M.Shabestari Target moved to empty
04/01/08 03:41 khem no rate on dE9R NEED TO BE CHECK
229270 04/01/08 03:35 A. Camsonne Rates run 2097
229269 04/01/08 03:35 A. Camsonne Rate check 1 uA C run 2097
229268 04/01/08 03:34 A. Camsonne Rate check 1 uA Carbon E DE
229266 04/01/08 03:25 M.Shabestari Target moved to Carbon
229261 04/01/08 02:52 Mitra Shabes Target moved to BeO
229254 04/01/08 02:11 A. Camsonne Hall A BCM rough calibration
229253 04/01/08 01:04 Moffit Missing reference channel events in TDC.
229252 04/01/08 00:46 jinhuang bigbite lib is updated
229251 04/01/08 00:43 V. Sulkosky Database Scripts added to L-HRS Start of Run
229250 04/01/08 00:35 Moffit BigBite trigger plane ADC/TDC histograms added to output
229248 04/01/08 00:08 Annand Swing shift summary